### v1

# Is Sarcasm Detection A Step-by-Step Reasoning Process in Large Language Models?

Abstract
Elaborating a series of intermediate reasoning steps significantly improves the ability of large language models (LLMs) to solve complex problems, as such steps would evoke LLMs to think sequentially. However, human sarcasm understanding is often considered an intuitive and holistic cognitive process, in which various linguistic, contextual, and emotional cues are integrated to form a comprehensive understanding of the speaker’s true intention, which is argued not be limited to a step-by-step reasoning process. To verify this argument, we introduce a new prompting framework called SarcasmCue, which contains four prompting strategies, $viz.$ chain of contradiction (CoC), graph of cues (GoC), bagging of cues (BoC) and tensor of cues (ToC), which elicits LLMs to detect human sarcasm by considering sequential and non-sequential prompting methods. Through a comprehensive empirical comparison on four benchmarking datasets, we show that the proposed four prompting methods outperforms standard IO prompting, CoT and ToT with a considerable margin, and non-sequential prompting generally outperforms sequential prompting.
1 Introduction
Figure 1: The comparison of the processes of mathematical reasoning and sarcasm detection.
Sarcasm is a subtle linguistic phenomenon that uses rhetorical devices such as hyperbole and figuration to convey true sentiments and intentions that are opposite to the literal meanings of the words used Wen et al. (2023); Zhang et al. (2023b). Sarcasm detection aims to combine different types of cues, such as linguistic features, contextual information, emotional knowledge, to form a comprehensive understanding of the author’s sarcastic attitude. Owing to its inherent ambivalence and figurative nature, sarcasm detection has persistently proven a formidable challenge spanning the eras from feature engineering to prompt engineering Yue et al. (2023); Zhang et al. (2023a).
Recent large language models have demonstrated impressive performance in downstream natural language processing (NLP) tasks, in which “System 1” - the fast, unconscious, and intuitive tasks, e.g., sentiment classification, topic analysis, etc., have been argued to be successfully performed Cui et al. (2024). Instead, increasing efforts have been devoted to the other class of tasks - “System 2”, which requires slow, deliberative and multi-step thinking, such as logical, mathematical, and commonsense reasoning tasks Wei et al. (2022). To improve the ability of LLMs to solve such complex problems, a widely adopted technique is to decompose complex problems into a series of intermediate solution steps prior to answer generation, and elicit LLMs to think step-by-step, such as chain of thought (CoT) Wei et al. (2022), tree of thought (ToT) Yao et al. (2024), graph of thought (GoT) Besta et al. (2024), etc.
However, sarcasm detection, as a holistic, intuitive, and non-rational cognitive process, is arguably in noncompliance with step-by-step logical reasoning due to two main reasons: (1) sarcasm expression does not strictly conform to formal logical structures, such as the law of hypothetical syllogism (i.e., if $\mathcal{A}\Rightarrow\mathcal{B}$ and $\mathcal{B}\Rightarrow\mathcal{C}$, then $\mathcal{A}\Rightarrow\mathcal{C}$). For example, “Poor Alice has fallen for that stupid Bob; and that stupid Bob is head over heels for Claire; but don’t assume for a second that Alice would like Claire”; (2) sarcasm judgment is typically a fluid combination of various cues, where each cue holds equal importance to the judgment of sarcasm, and there is no rigid sequence of steps among them. As shown in Fig. 1, linguistic, contextual and emotional factors are all crucial for rendering the sentence as sarcastic. Hence, the main research question can be summarized as:
RQ: Is human sarcasm detection a step-by-step reasoning process?
To answer this question, we propose a theoretical framework, called SarcasmCue, based on the sequential and non-sequential prompting paradigm. It consists of four prompting methods, i.e., chain of contradiction (CoC), graph of cues (GoC), bagging of cues (BoC) and tensor of cues (ToC). A cue is similar to a thought, which is concretely a coherent language sequence related to linguistics, context, or emotion that serves as an intermediate indicator toward identifying sarcasm, such as rhetorical devices, emotional words, etc. Each of the four prompting methods has its own focus and advantages. Specifically,
  • CoC. It builds upon CoT prompting and harnesses the quintessential property of sarcasm (namely the contradiction between surface sentiment and true intention). It aims to: (1) identify the literal meaning and surface sentiment by extracting keywords, sentimental phrases, etc.; (2) deduce the true intention by scrutinizing special punctuation, rhetorical devices, cultural background, etc.; and (3) determine the inconsistency between surface sentiment and true intention. It has a typical linear structure.
  • GoC. Generalizing over CoC, GoC frames the problem of sarcasm detection as a search over a graph and treats various cues (e.g., linguistic, contextual, emotional cues, etc.) as nodes, with the relations across cues represented as edges. Different from CoC and ToT, It allows language models to flexibly choose and weigh multiple cues when detecting sarcasm, rather than following a fixed hierarchy or linear reasoning path, unconstrained by the need for unique predecessor nodes. It represents a graphical structure. In summary, both CoC and GoC follow a step-by-step reasoning process.
  • BoC. In contrast, BoC and ToC are proposed based on the assumption that sarcasm detection is not a step-by-step reasoning process. BoC is a bagging approach that constructs a pool of diverse cues and creates multiple cue subsets through randomly sampling q cues at each round. LLMs are employed to generate multiple predictions based on these subsets, and such predictions are aggregated to produce the final result via majority voting. It has a set-based structure.
  • ToC. ToC treats each type of cues (namely linguistic, contextual, and emotional cues) as an independent, orthogonal view for sarcasm understanding and constructs a multi-view representation through the tensor product of these three types of cues. It allows language models to leverage higher-order interactions among the cues. ToC can be visualized as a 3D volumetric structure, where each coordinate axis corresponds to a distinct type of cue. This tensorial method aims to offer a more comprehensive and expressive means of fusing diverse cues.
We present empirical evaluations of the proposed prompting approaches across four sarcasm detection benchmarks over 2 SOTA LLMs (i.e., GPT-4o, LLaMA 3-8B), and compare their results against 3 SOTA prompting approaches (i.e., standard IO prompting, CoT, ToT). We show that the proposed four prompting methods outperforms standard IO prompting, CoT and ToT with a margin of 2%, and non-sequential prompting generally outperforms sequential prompting. Between the two LLMs, GPT-4o consistently beats LLaMA by a striking margin across all tasks.
The main contributions are concluded as follows:
  • Our work is the first to investigate the step-wise nature of sarcasm judgment by using both sequential and non-sequential prompting methods.
  • We propose a new prompting framework that consists of four sub-methods, $viz.$ chain of contradiction (CoC), graph of cues (GoC), bagging of cues (BoC) and tensor of cues (ToC).
  • Comprehensive experiments over four datasets demonstrate the superiority of the proposed prompting framework in zero-shot sarcasm detection.
2 Related Work
This section reviews two lines of research that form the basis of this work: CoT prompting and sarcasm detection.
2.1 Chain-of-Thought Prompting
Inspired by the step-by-step thinking ability of humans, CoT prompting was proposed to “prompt” language models to produce intermediate reasoning steps that lead to the final answer. Wei et al. Wei et al. (2022) made a formal definition of CoT prompting in LLMs and proved its effectiveness by presenting empirical evaluations on arithmetic reasoning benchmarks. This work pioneered the use of CoT prompting in NLP. However, its performance hinged on the quality of manually crafted prompts, which was a costly and unstable process. To fill this gap, Auto-CoT was proposed to automatically construct demonstrations with questions and reasoning chains Zhang et al. (2022). Different from Auto-CoT, Diao et al. Diao et al. (2023) presented an Active-Prompt approach to determine which questions were the most important and helpful to annotate from a pool of task-specific queries, for reducing the human engineering workload. The impressive results of CoT prompting have sparked a surge of exploration into designing CoT prompting strategies across various tasks Fei et al. (2023); Li et al. (2023); Zheng et al. (2023). For instance, Wang et al. Wang et al. (2024) used formal grammars as the intermediate reasoning steps for domain-specific language generation.
Furthermore, Yao et al. Yao et al. (2024) introduced a non-chain prompting framework, namely ToT, which made LLMs consider multiple different reasoning paths and self-evaluated choices to decide the next course of action. They proved the effectiveness of the ToT approach on the tasks requiring non-trivial planning or search. Beyond CoT and ToT approaches, Besta et al. Besta et al. (2024) modeled the information generated by an LLM as an arbitrary graph (i.e., GoT), where units of information were considered as vertices and the dependencies between these vertices were edges. Although the above-mentioned approaches have shown exceptional performance on various arithmetic and logical reasoning tasks, all of them adopt the sequential decoding paradigm of “let LLMs think step by step”. Contrarily, it is argued that sarcasm judgment does not conform to step-by-step logical reasoning, and there is a need to develop non-sequential prompting approaches.
Figure 2: An illustration of our SarcasmCue framework that consists of four prompting sub-methods.
2.2 Sarcasm Detection
Sarcasm detection is habitually treated as a text classification task, where the target is to identify whether the given text is sarcastic or not Zhang et al. (2024). It has evolved from early rule based and statistical learning based approaches to traditional neural methods, such as CNN, RNN, and further advanced to modern neural methods epitomized by Transformer models. In early stage, the rule based approaches infer the overall sarcasm polarity based on the refined sarcasm rules, such as the occurrence of the interjection word Zhang et al. (2023a). Statistical learning based approaches mainly employ statistical learning techniques, e.g., SVM, RF, NB, etc., to extract patterns and relationships within the data Zhou et al. (2023).
As deep learning based architectures have shown the superiority over statistical learning, numerous base neural networks, e.g., such as CNN Jain et al. (2020), LSTM Ghosh et al. (2018), GCN Liang et al. (2022), etc., have been predominantly utilized during the middle stage of sarcasm detection research, aiming to learn and extract complex features in an end-to-end fashion. As the field of deep learning continues to evolve, sarcasm detection research has stepped into the era of pre-trained language models (PLMs). An increasing number of researchers are designing sophisticated PLM architectures to serve as encoders for obtaining effective text representations. For example, Liu et al. Liu et al. (2022) proposed a dual-channel framework by modeling both literal and implied sentiments separately. They also constructed two conflict prompts to elicit PLMs to generate the sarcasm polarity Liu et al. (2023b). Qiao et al. Qiao et al. (2023) presented a mutual-enhanced incongruity learning network to take advantage of the underlying consistency between the two modules to boost the performance. Tian et al. Tian et al. (2023) proposed a dynamic routing Transformer network to activate different routing transformer modules for modeling the dynamic mechanism in sarcasm detection.
However, the above-mentioned works still focus on how to utilize PLMs to extract effective features, without leveraging the extraodinary context learning capabilities of LLMs. In contrast, this paper makes the first attempt to explore the potential of prompting LLMs in sarcasm detection.
3 The Proposed Framework: SarcasmCue
The overall schematic illustration of the proposed SarcasmCue framework is illustrated in Fig. 2. We qualitatively compare SarcasmCue to other prompting approaches in Tab. 1. SarcasmCue is the only one to fully support chain-based, tree-based, graph-based, set-based and multidimensional array-based reasoning. It is also the only one that simultaneously supports both sequential and non-sequential prompting methods.
Table 1: Comparison of prompting methods.
Scheme Seq? Non-Seq?
Chain? Tree? Grap? Set? Tensor?
[…]
3.1 Task Definition
Consider a sarcasm detection task. Given the data set $\mathcal{D}=\left\{\left(\mathcal{X},\mathcal{Y}\right)\right\}$, where $\mathcal{X}=\{x_{1},x_{2},\ldots,x_{n}\}$ denotes the input text sequence and $\mathcal{Y}=\{y_{1},y_{2},\ldots,y_{n}\}$ denotes the output label sequence. We use $\mathcal{L}_{\theta}$ to represent a large language model with parameter $\theta$. Our task is to leverage a collection of cues $\mathcal{C}=\{c_{1},c_{2},...,c_{k}\}$ to bridge the input $\mathcal{X}$ and the output $\mathcal{Y}$, where each cue $c_{i}$ is a coherent language sequence related to linguistics, context, or emotion that serves as an intermediate indicator toward identifying sarcasm.
3.2 Chain of Contradiction
We capture the inherent paradoxical nature of sarcasm, which is the incongruity between the surface sentiment and the true intention, and introduce chain of contradiction, a CoT-style paradigm that allows LLMs to decompose the problem of sarcasm detection into intermediate steps and solve each before making decision (Fig. 2 (a)). Each cue $c_{k}\sim\mathcal{L}_{\theta}^{CoC}\left(c_{k}|\mathcal{X},c_{1},c_{2},...,c_{k-1}\right)$ is sampled sequentially, then the output $\mathcal{Y}\sim\mathcal{L}_{\theta}^{CoC}\left(\mathcal{Y}|\mathcal{X},c_{1},...,c_{k}\right)$. A specific instantiation of CoC involves three steps:
Step 1. We first ask LLM to detect the surface sentiment via the following prompt $p_{1}$:
Given the input sentence [$\mathcal{X}$], what is the SURFACE sentiment, as indicated by clues such as keywords, sentimental phrases, emojis?
The output sequence $y_{1}\sim\mathcal{L}_{\theta}^{CoC}\left(\mathcal{Y}|p_{1}\right)$ is generated from the language model $\mathcal{L}_{\theta}^{CoC}$ conditioned on input prompt $p_{1}$.
Step 2. We then ask LLM to carefully discover the true intention via the following prompt $p_{2}$:
Deduce what the sentence really means, namely the TRUE intention, by carefully checking any rhetorical devices, language style, unusual punctuation, common senses.
The output sequence, denoted as $y_{2}$, is generated from the language model conditioned on prompt $p_{2}$ as well as the previous interaction $p_{1},y_{1}$, formulated as $y_{2}\sim\mathcal{L}_{\theta}^{CoC}\left(\mathcal{Y}|p_{1},y_{1},p_{2}\right).$
Step 3. We finally ask LLM to examine the consistency between surface sentiment and true intention and make the final prediction:
Based on Step 1 and Step 2, evaluate whether the surface sentiment aligns with the true intention. If they do not match, the sentence is probably ‘Sarcastic’. Otherwise, the sentence is ‘Not Sarcastic’. Return the label only.
$y_{3}$ is therefore generated based on a joint understanding of the preceding context $y_{1}$, $y_{2}$ and $p_{1}$, $p_{2}$, $p_{3}$: $y_{3}\sim\mathcal{L}_{\theta}^{CoC}\left(\mathcal{Y}|p_{1},y_{1},p_{2},y_{2},p_{3}\right)$. The sarcasm label is identified from $y_{3}$ as the output of CoC.
Notably, CoC is built based on the presumption that all the cues are linearly correlated, and detects human sarcasm through step-by-step reasoning. Different from the original CoT, however, the steps are explicitly designed for the sarcasm detection context. Further details are presented in Algorithm 1 in App. A.
3.3 Graph of Cues
The linear structure of CoC restricts it to a single path of reasoning. To fill this gap, we introduce graph of cues, a GoT-style paradigm that allows LLMs to flexibly choose and weigh multiple cues, unconstrained by the need for unique predecessor nodes (Fig. 2 (b)). GoC frames the problem of sarcasm detection as a search over a graph, and is formulated as a tuple $\left(\mathcal{M},\mathcal{G},\mathcal{E}\right)$, where $\mathcal{M}$ is the cue maker used to define what are the common cues, $\mathcal{G}$ is a graph of “sarcasm detection process”, $\mathcal{E}$ is cue evaluator used to determine which cues to keep selecting and in which order. Unlike ToT and GoT, GoC does not involve the modules of “thought generator” and “thought aggregation”.
1. Cue maker. Human sarcasm judgment often relies on the combination and analysis of one or more cues to achieve an accurate understanding. Such cues can be broadly categorized into three types: linguistic cues, contextual cues and emotional cues. Linguistic cues refer to the linguistic features inherent in the text, including keywords, rhetorical devices, punctuation and language style. Contextual cues refer to the environment and background of the text, including topic, cultural background, common knowledge. Emotional cues denote the emotional stance conveyed by the text, including emotional words, special symbols (such as emojis) and emotional contrasts. A total number of 4+3+3=10 cues are adopted.
2. Graph construction. In $\mathcal{G}=\left(V,E\right)$, the cues are regarded as vertices constituting the vertex set $V$, while the relations across cues form the edge set $E$. If there is an edge between cues $c_{k}$ and $c_{j}$, it is considered that $c_{k}$ and $c_{j}$ are closely related. Given the cue $c_{k}$, the cue evaluator $\mathcal{E}$ considers cue $c_{j}$ to provide the most complementary information to $c_{k}$, which would combine with $c_{k}$ to facilitate a deep understanding of sarcasm.
3. Cue evaluator. We involve $\mathcal{G}$ in the LLM detecting sarcasm process. To advance this process, the cue evaluator $\mathcal{E}$ assesses the current progress towards judging sarcasm by means of determining whether the cumulative cues obtained so far are sufficient to yield an accurate judgment. If so, the search goes to an end. Otherwise, it serves as a heuristic for the search algorithm, determining which additional cues to select and in what order, to further the detection process. Similar to ToT, an LLM is used as the cue evaluator $\mathcal{E}$.
We employ a voting strategy to determine the most valuable cue for selection, by explicitly comparing multiple potential cue candidates in a voting prompt, such as:
Given an input text $\mathcal{X}$, the target is to accurately detect sarcasm. Now, we have collected the keyword information as the first step: {keywords}, judge if this provides over 95% confidence for accurate detection. If so, output the result. Otherwise, from the remaining cues {rhetorical devices, punctuation, …}, vote the most valuable one to improve accuracy and confidence for the next step. This step can be formulated as $\mathcal{E}\left(\mathcal{L}_{\theta}^{GoC},c_{j+1}\right)\sim Vote\left\{\mathcal{L}_{\theta}^{GoC}\left(c_{j+1}|\mathcal{X},c_{1,2,...,j}\right)\right\}_{c_{j+1}\in\{c_{j+1},...,c_{k}\}}$. In a nutshell, it greedily selects the most valuable cue until the final judgment is reached.
Although the GoC enables the exploration of many possible paths across the cue graph, its nature remains grounded in a step-by-step reasoning paradigm (see Algorithm 2 in App. A).
3.4 Bagging of Cues
We further relax the assumption that the cues for sarcasm detection are inter-related. We introduce bagging of cues, an ensemble learning based paradigm that allows LLMs to independently consider varied combinations of cues without assuming a fixed order or dependency among them (Fig. 2 (c)).
BoC constructs a pool of the pre-defined $k=10$ cues $\mathcal{C}$. From this pool, $\mathcal{T}$ subsets are random sampled, each consisting of $q~{}\left(i.e.,1\leq q\leq k\right)$ cues. BoC thus leverages LLMs to generate $\mathcal{T}$ independent sarcasm predictions $\hat{y}_{t}$ based on the cues of each subset. Finally, such predictions are aggregated using a majority voting mechanism to produce the final sarcasm detection result. This approach embraces randomness in cue selection, enhancing the LLM’s ability to explore numerous potential paths, thus improving the robustness and accuracy of sarcasm detection. BoC consists of the following key steps:
Step 1. Cue subsets construction. A total of $\mathcal{T}$ cue subsets $\mathcal{S}_{t\in[1,2,...,\mathcal{T}]}=\left\{(c_{t_{1}},c_{t_{2}},...,c_{t_{q}}),t\in[1,2,...,\mathcal{T}]\right\}$ are created by randomly sampling without replacement from the complete pool of cues $\mathcal{C}$. Each sampling is independent.
Step 2. LLM prediction. For each subset $\mathcal{S}_{t}$, an LLM $\mathcal{L}_{\theta}^{BoC}$ is used to independently make sarcasm prediction through the comprehensive analysis of the cues in the subset and the input text. This can be conceptually encapsulated as $\hat{y}_{t}\sim\mathcal{L}_{\theta}^{BoC}\left(\mathcal{Y}|\mathcal{S}_{t},\mathcal{X}\right)$.
Step 3. Prediction aggregation. These individual predictions are then combined using an aggregation function, i.e., majority voting, to yield the final prediction: $Y\sim Vote\left(\left\{\hat{y}_{1},\hat{y}_{2},...,\hat{y}_{\mathcal{T}}\right\}\right)$.
BoC treats all cues as independent and does not follow the step-by-step reasoning paradigm for sarcasm detection (see Algorithm 3 in App. A).
3.5 Tensor of Cues
CoC and GoC methods mainly handle low-order interactions between cues, while BoC assumes cues are independent. To capture high-order interactions among cues, we introduce tensor of cues, a novel paradigm that allows LLMs to amalgamate three types of cues ($viz.$ liguistic, contextual and emotional cues) into a high-dimensional representation (Fig. 2 (d)).
ToC treats each type of cues as an independent, orthogonal view for sarcasm understanding, and constructs a multi-view representation through the tensor product of such three types of cues. We first ask the LLM to extract linguistic, contextual, and emotional cues respectively via a simple prompt. Taking linguistic cue extraction as an example:
Instruction: Please extract the linguistic cues from the input sentence for sarcasm detection, such as keywords, rhetorical devices, punctuation and language style. Input: [$\mathcal{X}$]
We take the outputs of the LLM’s final hidden layer as the embeddings of the linguistic, contextual and emotional cues, and apply a tensor fusion mechanism to fuse the cues as additional inputs to the sarcasm detection prompt. Inspired by the success of tensor fusion network (TFN) for multi-modal sentiment analysis Zadeh et al. (2017), we apply token-wise tensor fusion to aggregate the cues. In particular, the embeddings are projected on a low-dimensional space, i.e., $\vec{Lin}=\left(e_{1}^{l},e_{2}^{l},...,e_{L}^{l}\right)^{T}$, $\vec{Con}=\left(e_{1}^{c},e_{2}^{c},...,e_{L}^{c}\right)^{T}$, $\vec{Emo}=\left(e_{1}^{e},e_{2}^{e},...,e_{L}^{e}\right)^{T}$. Suppose the LLM has a hidden dimensionality of $d$, fully-connected layers $f_{lin},f_{con},f_{emo}$ are constructed to map the embeddings to dimensionality of $\{d_{l},d_{c},d_{e}\}$, respectively for linguistic, contextual and emotional cues. Then, a tensor product is computed to combine the cues into a high-dimensional representation $\mathcal{Z}=\left(e_{1},e_{2},...,e_{L}\right)^{T}$, where
$e_{i}=\begin{bmatrix}e_{i}^{l}\\ 1\end{bmatrix}\otimes\begin{bmatrix}e_{i}^{c}\\ 1\end{bmatrix}\otimes\begin{bmatrix}e_{i}^{e}\\ 1\end{bmatrix},\forall i\in[1,2,...,L].$ (1)
The additional value of 1 facilitates an explicit rendering of single-cue features and bi-cue interactions, leading to a comprehensive fusion of different cues encapsulated in each fused token $e_{i}\in\mathcal{R}^{(d_{l}+1)\times(d_{c}+1)\times(d_{e}+1)}$. The values of $d_{l}$, $d_{c}$ and $d_{e}$ are delicately chosen such that the dimensionality of fused token is precisely $d$ [^1] Otherwise the fused tokens are truncated to d-dim vectors. That enables an integration of the aggregated cues to the main prompt via:
Consider the information provided in the current cue above. Classify whether the input text is sarcastic or not. If you think the Input text is sarcastic, answer: yes. If you think the Input text is not sarcastic, answer: no. Input: [$\mathcal{X}$]
The embedded prompt above is prepended with the aggregated cue sequence $\mathcal{Z}$ before fed to the LLM. As it is expected to output a single token of “yes” or “no” by design, we take the logit of the first generated token and decode the label accordingly as the output of ToC.
ToC facilitates deep interactions among these cues, providing a powerful and flexible framework for processing complex linguistic phenomena (see Algorithm 4 in App. A). Notably, as ToC manipulates cues on the vector level via neural structures, it requires access to the LLM structure and calls for supervised training on a collection of labeled samples. During training, the weights of the LLM are frozen, and the linear weights in $f_{lin},f_{con},f_{emo}$ are updated as an adaptation of LLM to the task context.
4 Experiments
4.1 Experiment Setups
Datasets. Four benchmarking datasets are selected as the experimental beds, $viz.$ IAC-V1 Lukin and Walker (2013), IAC-V2 Oraby et al. (2016), SemEval 2018 Task 3 Van Hee et al. (2018) and MUStARD Castro et al. (2019).
IAC-V1 and IAC-V2 are from the Internet Argument Corpus (IAC) Lukin and Walker (2013), specifically designed for the task of identifying and analyzing sarcastic remarks within online debates and discussions. It encompasses a balanced mixture of sarcastic and non-sarcastic comments.
SemEval 2018 Task 3 is collected using irony-related hashtags (i.e. #irony, #sarcasm, #not) and are subsequently manually annotated to minimise the amount of noise in the corpuses. It emphasize the challenges inherent in identifying sarcasm within the constraints of MUStARD’s concise format, and highlight the importance of context and linguistic subtleties in recognizing sarcasm.
MUStARD is compiled from popular TV shows including Friends, The Golden Girls, The Big Bang Theory, etc. It consists of 690 samples total of 3,000 utterances. Each sample is a conversation consisting of several utterances. In this work, we only use the textual information. The statistics for each dataset are shown in Table 2.
Table 2: Dataset statistics.
Dataset Avg. Length #Train #Dev #Test
IAC-V1 68 1,595 80 320
IAC-V2 43 5,216 262 1,042
SemEval 2018 14 3,634 200 784
MUStARD 14 552 - 138
Baselines. A wide range of SOTA baselines are included for comparison. They are:
  • PLMs.  (1) RoBERTa Liu et al. (2019), (2) BNS-Net Zhou et al. (2023), (3) DC-Net Liu et al. (2022), (4) QUIET Liu et al. (2023a) and (5) SarcPrompt Liu et al. (2023b) are five SOTA PLMs based approaches for sarcasm detection via pre-trained language modeling and refined representations.
  • Prompt tuning.  (6) IO, (7) CoT Wei et al. (2022) and (8) ToT Yao et al. (2024) are four SOTA prompting approaches by leveraging advanced prompt approaches to enhance LLM’s performance.
  • LLMs.  (9) GPT-4o [^2] https://openai.com/index/hello-gpt-4o/ and (10) LLAMA 3-8B-Instruct [^3] https://llama.meta.com/llama3/ are the strongest general LLMs.
Implementation. We have implemented the prompting methods for GPT-4o and LLaMA3-8B-Instruct, and reported the performance of PLMs in their original papers. The GPT-4o methods are implemented with the official openAI Python API library [^4] https://github.com/openai/openai-python, while the LLaMA methods are implemented based on the Hugging Face Transformers library [^5] https://huggingface.co/docs/transformers. All prompting strategies are implemented for GPT-4o and LLaMA3-8B-Instruct except for ToC, which can solely be deployed on open-sourced LLMs. Following previous works in this field, LangChain [^6] https://github.com/langchain-ai/langchain is employed for the implementation of ToT and GoC. For the training of ToC, cross-entropy loss between the output logit and the true label token is computed to update the weights of the fully-connected layers.
Table 3: Performance on four datasets. For LLMs, all strategies but ToC are based on a zero-shot setting.
Paradigm Method IAC-V1 IAC-V2 SemEval 2018 MUStARD Avg. of F1
Acc. Ma-F1 Acc. Ma-F1 Acc. Ma-F1 Acc. Ma-F1
PLMs RoBERTa 72.10 72.90 82.70 82.70 73.90 72.80 66.27 65.16 73.39
BNS-Net 66.13 65.95 75.93 75.92 73.47 73.44 - - 71.77
DC-Net 66.50 66.40 82.10 82.10 76.70 76.30 - - 74.93
QUIET - - - - - - 72.36 72.13 -
SarcPrompt 75.20 75.20 84.90 84.90 76.90 76.60 66.58 66.63 75.78
GPT-4o IO 70.63 70.05 73.03 71.99 64.03 63.17 67.24 65.79 68.14
CoT 61.56 58.49 58.83 56.42 58.92 51.99 58.11 55.76 55.67
ToT 71.56 71.17 70.63 69.07 63.90 63.02 69.00 68.27 67.46
[…]
GoC (Ours) 85.38 68.08 64.97 61.30 74.03 74.02 70.69 69.91 68.33
BoC (Ours) 68.75 67.36 71.35 69.39 62.12 61.85 69.42 68.45 66.79
LLaMA 3-8B-Instruct IO 55.94 46.40 54.70 43.74 49.36 44.46 54.64 44.99 44.90
CoT 56.25 47.28 54.22 42.96 49.36 44.55 54.20 44.86 44.91
ToT 52.50 48.98 55.95 53.05 50.64 48.63 54.35 50.56 50.31
[…]
GoC (Ours) 57.10 54.96 42.20 41.61 57.33 57.24 52.77 52.67 51.62
BoC (Ours) 62.50 59.28 62.57 58.11 59.82 58.40 59.71 56.70 58.12
ToC (Ours) 70.31 70.29 79.08 79.07 77.93 76.86 73.33 72.85 74.77
4.2 Main Results
We report both Accuracy and Macro-F1 results for SarcasmCue and baselines in a zero-shot setting in Table 3, except for ToC which requires supervised training for context adaption.
LLMs do not possess a unique advantage on sarcasm detection. Since sarcasm indicates the manifestation of sentiments and intentions opposite to the literal meaning of the texts, it usually violates logical reasoning pipelines that LLMs are known to excel at Wei et al. (2022). This is empirically validated in the experiment where LLMs are observed to have consistently lower performance over PLMs in terms of average F1 scores across the four datasets. This highlights the need to investigate prompting strategies for adapting LLMs for sarcasm detection, towards which this work has made the first attempt and achieved preliminary success.
Human sarcasm detection does not necessarily follow a step-by-step reasoning process. The comparison between sequential (CoT, CoC, GoC, ToT) and non-sequential (BoC, ToC) prompting strategies fails to provide clear empirical evidences on whether sarcasm detection follows a step-by-step reasoning process. Nevertheless, the results on LLaMA3-8B-Instruct are more indicative to GPT-4o, since the latter has a strong capacity on its own (IO) and does not significantly benefit from any prompting strategies on its top. On LLaMA3-8B-Instruct where in-context learning is necessary for sarcasm detection due to its poor IO performance, non-sequential approaches can apparently offer more benefits over sequential ones, with a remarkable margin consistently present on all four datasets. This seems to support our hypothesize that sarcasm has a non-sequential nature.
SarcasmCue successfully adapts LLMs to sarcasm detection. The proposed prompting strategies in the SarcasmCue framework achieve an overall superior performance to the baseline prompting methods and bring about accuracy increase over the original LLMs in a zero-shot setting. In particular, by explicitly designing the reasoning steps for sarcasm detection, CoC beats CoT by a tremendous margin on GPT-4o, whilst performing in par with CoT on LLaMA3-8B-Instruct, an interesting result that further suggests the non-squential nature of sarcasm detection. By pre-defining the set of cues on 3 main aspects, GoC and BoC manage to guide LLMs to reason along the correct paths, leading to more accuracy judgment of sarcasm than the freestyle thinking in ToT. The proposed trainable neural architecture in ToC achieves an effective tensor fusion of multi-aspect cues for sarcasm detection, pushing the capacity to a comparable level to PLMs without tuning the LLM parameters.
5 Acknowledgments
This work is supported by National Science Foundation of China under grant No. 62006212, Fellowship from the China Postdoctoral Science Foundation (2023M733907), Natural Science Foundation of Hunan Province of China (242300421412).
6 Conclusion
In this work, we aim to study the step-wise reasoning nature of sarcasm detection, and introduce a prompting framework (called SarcasmCue) containing four sub-methods, $viz.$ chain of contradiction (CoC), graph of cues (GoC), bagging of cues (BoC) and tensor of cues (ToC). It elicits LLMs for human sarcasm detection by considering sequential and non-sequential prompting methods. Our comprehensive evaluations across multiple benchmarks and state-of-the-art LLMs demonstrate that SarcasmCue outperforms traditional methods, with non-sequential prompting methods (GoC and ToC) showing particularly strong performance. In the future, we plan to develop the multi-modal version of SarcasmCue for multi-modal sarcasm detection.
7 Limitations
The proposed SarcasmCue model has several limitations: (1) It incorporates only three types of cues – linguistic, contextual, and emotional – while other potentially useful cues, such as multimodal information, have not been integrated, potentially limiting the model’s comprehensive understanding of sarcasm; (2) the performance of SarcasmCue is influenced by the capabilities of the underlying large language models (LLMs), meaning it performs better with more powerful LLMs.
Appendix A Algorithms of Four Prompting Methods
1. CoC. We present further details of CoC in Algorithm 1.
Algorithm 1 Chain of contradiction
1:
2:Input: Sentence $\mathcal{X}$, an LLM $\mathcal{L}_{\theta}$
3:
4:Output: Sarcasm Label $\mathcal{Y}$
5:Step 1: Detect surface sentiment
6:Output cue $c_{1}$: $c_{1}\sim\mathcal{L}_{\theta}^{CoC}\left(c_{1}|\mathcal{X},p_{1}\right)$
7:Step 2: Discover true intention
8:Output cue $c_{2}$: $c_{2}\sim\mathcal{L}_{\theta}^{CoC}\left(c_{2}|\mathcal{X},c_{1},p_{2}\right)$
9:Step 3: Evaluate consistency and make prediction
10:Output cue $c_{3}$: $c_{3}\sim\mathcal{L}_{\theta}^{CoC}\left(c_{3}|\mathcal{X},c_{1},c_{2},p_{3}\right)$
11:$\mathcal{Y}=\begin{cases}\text{Sarcastic}&\text{if }c_{1}\neq c_{2}\\ \text{Not Sarcastic}&\text{otherwise}\end{cases}$
12:return $\mathcal{Y}$
2. GoC. We present further details of GoC in Algorithm 2.
Algorithm 2 Graph of Cues (GoC) for Sarcasm Detection
1:
2:Input: Sentence $\mathcal{X}$, an LLM $\mathcal{L}_{\theta}$
3:
4:Output: Sarcasm Label $\mathcal{Y}$
5:1. Graph Construction
6:Construct graph $\mathcal{G}=(V,E)$ where 10 cues are vertices $V$ and relationships between cues are edges $E$
7:2. Sarcasm Detection Process
8:Initialize selected cues $C_{\text{selected}}=\emptyset$, $j=0$
9:Initialize current confidence $\mathbb{C}=0$
10:while $\mathbb{C}<0.95\cap j\leq 10$ do
11:    Select the most valuable cue:
12:    $c_{\text{j+1}}\sim Vote\left\{\mathcal{L}_{\theta}^{GoC}\left(c_{j+1}|\mathcal{X},c_{1},c_{2},...,c_{j}\right)\right\}_{c_{j+1}\in\{c_{j+1},...,c_{10}\}}$
13:    Add $c_{\text{j+1}}$ to $C_{\text{selected}}$
14:    Update current confidence $\mathbb{C}$, $j$++
15:Make final judgment based on $C_{\text{selected}}$: $\mathcal{Y}=\mathcal{L}_{\theta}^{GoC}\left(\mathcal{Y}|\mathcal{X},C_{\text{selected}}\right)$
16:return $\mathcal{Y}$
Algorithm 3 Bagging of cues
1:
2:Input: Sentence $\mathcal{X}$, Cue Pool $\mathcal{C}$, Number of Subsets $\mathcal{T}$, Number of Cues per Subset $q$, an LLM $\mathcal{L}_{\theta}$
3:
4:Output: Sarcasm Label $Y$
5:Step 1: Cue Subsets Construction
6:for $t=1$ to $\mathcal{T}$ do
7:    Randomly sample a subset $\mathcal{S}_{t}=\{c_{t1},c_{t2},\ldots,c_{tq}\}$ from $\mathcal{C}$
8:Step 2: LLM Prediction
9:for $t=1$ to $\mathcal{T}$ do
10:    Generate sarcasm prediction $\hat{y}_{t}\sim\mathcal{L}_{\theta}^{BoC}(\hat{y}_{t}|\mathcal{S}_{t},\mathcal{X})$
11:Step 3: Prediction Aggregation
12:Aggregate predictions using majority voting:
13:$Y\sim Vote(\{\hat{y}_{1},\hat{y}_{2},\ldots,\hat{y}_{\mathcal{T}}\})$
14:return $Y$
3. BoC. We present further details of BoC in Algorithm 3.
4. ToC. We present further details of ToC in Algorithm 4.
Algorithm 4 Tensor of cues
1:
2:Input: Sentence $\mathcal{X}$, an LLM $\mathcal{L}_{\theta}$
3:
4:Output: Sarcasm Label $\mathcal{Y}$
5:Step 1: Extract Cues
6:Obtain linguistic cue embeddings $\vec{Lin}=(e_{1}^{l},e_{2}^{l},\ldots,e_{m}^{l})^{T}$, contextual cue embeddings $\vec{Con}=(e_{1}^{c},e_{2}^{c},\ldots,e_{p}^{c})^{T}$, emotional cue embeddings $\vec{Emo}=(e_{1}^{e},e_{2}^{e},\ldots,e_{s}^{e})^{T}$
7:Step 2: Construct Tensor Representation
8:Compute tensor product to combine cues: $\mathcal{Z}=\vec{Lin}\otimes\vec{Con}\otimes\vec{Emo}$
9:Step 3: Sarcasm Detection
10:Take tensor $\mathcal{Z}$ as input to a LLM for sarcasm detection:
11:$\mathcal{Y}\sim\mathcal{L}_{\theta}^{ToC}(\mathcal{Y}|\mathcal{Z},\mathcal{X})$
12:return $\mathcal{Y}$

### v2

@@
 Elaborating a series of intermediate reasoning steps significantly improves the ability of large language models (LLMs) to solve complex problems, as such steps would evoke LLMs to think sequentially.
-However, human sarcasm understanding is often considered an intuitive and holistic cognitive process, in which various linguistic, contextual, and emotional cues are integrated to form a comprehensive understanding of the speaker’s true intention, which is argued not be limited to a step-by-step reasoning process.
-To verify this argument, we introduce a new prompting framework called SarcasmCue, which contains four prompting strategies, $viz.$ chain of contradiction (CoC), graph of cues (GoC), bagging of cues (BoC) and tensor of cues (ToC), which elicits LLMs to detect human sarcasm by considering sequential and non-sequential prompting methods.
-Through a comprehensive empirical comparison on four benchmarking datasets, we show that the proposed four prompting methods outperforms standard IO prompting, CoT and ToT with a considerable margin, and non-sequential prompting generally outperforms sequential prompting.
+However, human sarcasm understanding is often considered an intuitive and holistic cognitive process, in which various linguistic, contextual, and emotional cues are integrated to form a comprehensive understanding, in a way that does not necessarily follow a step-by-step fashion.
+To verify the validity of this argument, we introduce a new prompting framework (called SarcasmCue) containing four sub-methods, $viz.$ chain of contradiction (CoC), graph of cues (GoC), bagging of cues (BoC) and tensor of cues (ToC), which elicits LLMs to detect human sarcasm by considering sequential and non-sequential prompting methods.
+Through a comprehensive empirical comparison on four benchmarks, we highlight three key findings: (1) CoC and GoC show superior performance with more advanced models like GPT-4 and Claude 3.5, with an improvement of 3.5% $\uparrow$.
+(2) ToC significantly outperforms other methods when smaller LLMs are evaluated, boosting the F1 score by 29.7% $\uparrow$ over the best baseline.
+(3) Our proposed framework consistently pushes the state-of-the-art (i.e., ToT) by 4.2%, 2.0%, 29.7%, and 58.2% in F1 scores across four datasets.
+This demonstrates the effectiveness and stability of the proposed framework.
+[^1] Our codes are available at https://github.com/qiuchili/llm_sarcasm_detection.
 1 Introduction
 Figure 1: The comparison of the processes of mathematical reasoning and sarcasm detection.
-Sarcasm is a subtle linguistic phenomenon that uses rhetorical devices such as hyperbole and figuration to convey true sentiments and intentions that are opposite to the literal meanings of the words used Wen et al.
-(2023); Zhang et al.
-(2023b).
-Sarcasm detection aims to combine different types of cues, such as linguistic features, contextual information, emotional knowledge, to form a comprehensive understanding of the author’s sarcastic attitude.
-Owing to its inherent ambivalence and figurative nature, sarcasm detection has persistently proven a formidable challenge spanning the eras from feature engineering to prompt engineering Yue et al.
-(2023); Zhang et al.
-(2023a).
-Recent large language models have demonstrated impressive performance in downstream natural language processing (NLP) tasks, in which “System 1” - the fast, unconscious, and intuitive tasks, e.g., sentiment classification, topic analysis, etc., have been argued to be successfully performed Cui et al.
+Recent large language models have demonstrated impressive performance across downstream natural language processing (NLP) tasks, in which “System 1” - the fast, unconscious, and intuitive tasks, e.g., sentiment classification, topic analysis, etc., have been argued to be successfully performed Cui et al.
 (2024).
-Instead, increasing efforts have been devoted to the other class of tasks - “System 2”, which requires slow, deliberative and multi-step thinking, such as logical, mathematical, and commonsense reasoning tasks Wei et al.
+Instead, increasing efforts have been devoted to the other class of tasks - “System 2”, which requires slow, deliberative and multi-steps thinking, such as logical, mathematical, and commonsense reasoning tasks Wei et al.
 (2022).
-To improve the ability of LLMs to solve such complex problems, a widely adopted technique is to decompose complex problems into a series of intermediate solution steps prior to answer generation, and elicit LLMs to think step-by-step, such as chain of thought (CoT) Wei et al.
+To improve the ability of LLMs to solve such complex problems, a popular paradigm is to decompose complex problems into a series of intermediate solution steps, and elicit LLMs to think step-by-step, such as chain of thought (CoT) Wei et al.
 (2022), tree of thought (ToT) Yao et al.
@@
 (2024), etc.
-However, sarcasm detection, as a holistic, intuitive, and non-rational cognitive process, is arguably in noncompliance with step-by-step logical reasoning due to two main reasons: (1) sarcasm expression does not strictly conform to formal logical structures, such as the law of hypothetical syllogism (i.e., if $\mathcal{A}\Rightarrow\mathcal{B}$ and $\mathcal{B}\Rightarrow\mathcal{C}$, then $\mathcal{A}\Rightarrow\mathcal{C}$).
-For example, “Poor Alice has fallen for that stupid Bob; and that stupid Bob is head over heels for Claire; but don’t assume for a second that Alice would like Claire”; (2) sarcasm judgment is typically a fluid combination of various cues, where each cue holds equal importance to the judgment of sarcasm, and there is no rigid sequence of steps among them.
-As shown in Fig.
-1, linguistic, contextual and emotional factors are all crucial for rendering the sentence as sarcastic.
+However, due to its inherent ambivalence and figurative nature, sarcasm detection is often considered a holistic and non-rational cognitive process that does not conform to step-by-step logical reasoning for two main reasons: (1) sarcasm expression does not strictly conform to formal logical structures, such as the law of hypothetical syllogism (i.e., if $\mathcal{A}\Rightarrow\mathcal{B}$ and $\mathcal{B}\Rightarrow\mathcal{C}$, then $\mathcal{A}\Rightarrow\mathcal{C}$).
+For example, “Poor Alice has fallen for that stupid Bob; and that stupid Bob is head over heels for Claire; but don’t assume for a second that Alice would like Claire”; (2) sarcasm judgment is often considered a fluid combination of various cues.
+Each cue holds equal importance and there is no rigid sequence of steps among them, as shown in Fig.
+1.
 Hence, the main research question can be summarized as:
@@
 It consists of four prompting methods, i.e., chain of contradiction (CoC), graph of cues (GoC), bagging of cues (BoC) and tensor of cues (ToC).
-A cue is similar to a thought, which is concretely a coherent language sequence related to linguistics, context, or emotion that serves as an intermediate indicator toward identifying sarcasm, such as rhetorical devices, emotional words, etc.
-Each of the four prompting methods has its own focus and advantages.
-Specifically,
+Each method has its own focus and advantages.
+In this work, cue is similar to thought, being a coherent language sequence related to linguistics, context, or emotion that serves as an intermediate indicator for identifying sarcasm, such as rhetorical devices or emotional words.
+More specifically,
   • CoC.
-It builds upon CoT prompting and harnesses the quintessential property of sarcasm (namely the contradiction between surface sentiment and true intention).
-It aims to: (1) identify the literal meaning and surface sentiment by extracting keywords, sentimental phrases, etc.; (2) deduce the true intention by scrutinizing special punctuation, rhetorical devices, cultural background, etc.; and (3) determine the inconsistency between surface sentiment and true intention.
-It has a typical linear structure.
+It harnesses the quintessential property of sarcasm (namely the contradiction between surface sentiment and true intention).
+It aims to: (1) identify the surface sentiment by extracting keywords, etc.; (2) deduce the true intention by scrutinizing rhetorical devices, etc.; and (3) determine the inconsistency between them.
+It is a typical linear structure.
   • GoC.
-Generalizing over CoC, GoC frames the problem of sarcasm detection as a search over a graph and treats various cues (e.g., linguistic, contextual, emotional cues, etc.) as nodes, with the relations across cues represented as edges.
-Different from CoC and ToT, It allows language models to flexibly choose and weigh multiple cues when detecting sarcasm, rather than following a fixed hierarchy or linear reasoning path, unconstrained by the need for unique predecessor nodes.
-It represents a graphical structure.
-In summary, both CoC and GoC follow a step-by-step reasoning process.
+Generalizing over CoC, GoC frames the problem of sarcasm detection as a search over a graph and treats various cues as nodes, with the relations across cues represented as edges.
+Unlike CoC and ToT, it goes beyond following a fixed hierarchy or linear reasoning path.
+In summary, both CoC and GoC follow the step-by-step reasoning process.
   • BoC.
-In contrast, BoC and ToC are proposed based on the assumption that sarcasm detection is not a step-by-step reasoning process.
-BoC is a bagging approach that constructs a pool of diverse cues and creates multiple cue subsets through randomly sampling q cues at each round.
-LLMs are employed to generate multiple predictions based on these subsets, and such predictions are aggregated to produce the final result via majority voting.
-It has a set-based structure.
+BoC is a bagging approach that constructs a pool of diverse cues and randomly sampling multiple cue subsets.
+LLMs are employed to generate multiple predictions based on these subsets, and such predictions are aggregated to produce the final result.
+It is a set-based structure.
   • ToC.
-ToC treats each type of cues (namely linguistic, contextual, and emotional cues) as an independent, orthogonal view for sarcasm understanding and constructs a multi-view representation through the tensor product of these three types of cues.
+ToC treats each type of cues (namely linguistic, contextual, and emotional cues) as an independent, orthogonal view for sarcasm understanding and constructs a multi-view representation through the tensor product.
 It allows language models to leverage higher-order interactions among the cues.
-ToC can be visualized as a 3D volumetric structure, where each coordinate axis corresponds to a distinct type of cue.
-This tensorial method aims to offer a more comprehensive and expressive means of fusing diverse cues.
-We present empirical evaluations of the proposed prompting approaches across four sarcasm detection benchmarks over 2 SOTA LLMs (i.e., GPT-4o, LLaMA 3-8B), and compare their results against 3 SOTA prompting approaches (i.e., standard IO prompting, CoT, ToT).
-We show that the proposed four prompting methods outperforms standard IO prompting, CoT and ToT with a margin of 2%, and non-sequential prompting generally outperforms sequential prompting.
-Between the two LLMs, GPT-4o consistently beats LLaMA by a striking margin across all tasks.
+ToC can be visualized as a 3D volumetric structure.
+Hence, BoC and ToC are proposed based on the assumption that sarcasm detection is not a step-by-step reasoning process.
+  • Their correlation.
+These four methods represent an evolution from linear to nonlinear, and from a single perspective to multiple perspectives, together forming a comprehensive theoretical framework (SarcasmCue).
+Their design aims to adapt to various sarcasm detection scenarios.
+We present empirical evaluations of the proposed prompting approaches across four benchmarks over 4 SOTA LLMs (i.e., GPT-4o, Claude 3.5 Sonnet, Llama 3-8B, Qwen 2-7B), and compare their results against 3 SOTA prompting approaches (i.e., standard IO prompting, CoT and ToT).
+we highlight three key observations: (1) When the base model is more advanced (such as GPT-4 and Claude 3.5 Sonnet), CoC and GoC show superior performance against the state-of-the-art (SoTA) baseline with an improvement of 3.5% $\uparrow$.
+(2) ToC achieves the best performance when smaller LLMs are evaluated.
+For example, in Llama 3-8B, ToC’s average F1 score of 65.24 represents a 29.7% improvement over the best baseline method, ToT.
+In Qwen 2-7B, ToC shows a 58.2% improvement over the best baseline method, IO.
+(3) Our proposed framework consistently pushes SoTA by 4.2%, 2.0%, 29.7% and 58.2% in F1 scores across four datasets.
+This demonstrates the effectiveness of the proposed framework.
 The main contributions are concluded as follows:
-  • Our work is the first to investigate the step-wise nature of sarcasm judgment by using both sequential and non-sequential prompting methods.
-  • We propose a new prompting framework that consists of four sub-methods, $viz.$ chain of contradiction (CoC), graph of cues (GoC), bagging of cues (BoC) and tensor of cues (ToC).
-  • Comprehensive experiments over four datasets demonstrate the superiority of the proposed prompting framework in zero-shot sarcasm detection.
+  • Our work is the first to investigate the stepwise reasoning nature of sarcasm detection by using both sequential and non-sequential prompting methods.
+  • We propose a new prompting framework that consists of four sub-methods, $viz.$ CoC, GoC, BoC and ToC.
+  • Comprehensive experiments over four datasets demonstrate the superiority of the proposed prompting framework.
 2 Related Work
-This section reviews two lines of research that form the basis of this work: CoT prompting and sarcasm detection.
 2.1 Chain-of-Thought Prompting
-Inspired by the step-by-step thinking ability of humans, CoT prompting was proposed to “prompt” language models to produce intermediate reasoning steps that lead to the final answer.
+Inspired by the step-by-step thinking ability of humans, CoT prompting was proposed to “prompt” language models to produce intermediate reasoning steps.
 Wei et al.
@@
 (2022) made a formal definition of CoT prompting in LLMs and proved its effectiveness by presenting empirical evaluations on arithmetic reasoning benchmarks.
-This work pioneered the use of CoT prompting in NLP.
-However, its performance hinged on the quality of manually crafted prompts, which was a costly and unstable process.
+However, its performance hinged on the quality of manually crafted prompts.
 To fill this gap, Auto-CoT was proposed to automatically construct demonstrations with questions and reasoning chains Zhang et al.
 (2022).
-Different from Auto-CoT, Diao et al.
-Diao et al.
-(2023) presented an Active-Prompt approach to determine which questions were the most important and helpful to annotate from a pool of task-specific queries, for reducing the human engineering workload.
-The impressive results of CoT prompting have sparked a surge of exploration into designing CoT prompting strategies across various tasks Fei et al.
-(2023); Li et al.
-(2023); Zheng et al.
-(2023).
-For instance, Wang et al.
-Wang et al.
-(2024) used formal grammars as the intermediate reasoning steps for domain-specific language generation.
 Furthermore, Yao et al.
 Yao et al.
-(2024) introduced a non-chain prompting framework, namely ToT, which made LLMs consider multiple different reasoning paths and self-evaluated choices to decide the next course of action.
-They proved the effectiveness of the ToT approach on the tasks requiring non-trivial planning or search.
+(2024) introduced a non-chain prompting framework, namely ToT, which made LLMs consider multiple different reasoning paths to decide the next course of action.
 Beyond CoT and ToT approaches, Besta et al.
@@
 (2024) modeled the information generated by an LLM as an arbitrary graph (i.e., GoT), where units of information were considered as vertices and the dependencies between these vertices were edges.
-Although the above-mentioned approaches have shown exceptional performance on various arithmetic and logical reasoning tasks, all of them adopt the sequential decoding paradigm of “let LLMs think step by step”.
-Contrarily, it is argued that sarcasm judgment does not conform to step-by-step logical reasoning, and there is a need to develop non-sequential prompting approaches.
+However, all of them adopt the sequential decoding paradigm of “let LLMs think step by step”.
+Contrarily, it is argued that sarcasm judgment does not conform to step-by-step logical reasoning, and there is an urgent need to develop non-sequential prompting approaches.
 Figure 2: An illustration of our SarcasmCue framework that consists of four prompting sub-methods.
 2.2 Sarcasm Detection
-Sarcasm detection is habitually treated as a text classification task, where the target is to identify whether the given text is sarcastic or not Zhang et al.
+Sarcasm detection has evolved from early statistical learning based approaches to traditional neural methods, and further advanced to modern neural methods epitomized by Transformer models.
+In early stage, statistical learning based approaches mainly employ statistical learning techniques, e.g., SVM, NB, etc., to extract patterns and relationships within the data Zhang et al.
 (2023).
-It has evolved from early rule based and statistical learning based approaches to traditional neural methods, such as CNN, RNN, and further advanced to modern neural methods epitomized by Transformer models.
-In early stage, the rule based approaches infer the overall sarcasm polarity based on the refined sarcasm rules, such as the occurrence of the interjection word Zhang et al.
-(2023a).
-Statistical learning based approaches mainly employ statistical learning techniques, e.g., SVM, RF, NB, etc., to extract patterns and relationships within the data Zhou et al.
-(2023).
-As deep learning based architectures have shown the superiority over statistical learning, numerous base neural networks, e.g., such as CNN Jain et al.
+As deep learning based architectures have shown the superiority, numerous base neural networks, e.g., such as CNN Jain et al.
 (2020), LSTM Ghosh et al.
 (2018), GCN Liang et al.
-(2022), etc., have been predominantly utilized during the middle stage of sarcasm detection research, aiming to learn and extract complex features in an end-to-end fashion.
-As the field of deep learning continues to evolve, sarcasm detection research has stepped into the era of pre-trained language models (PLMs).
-An increasing number of researchers are designing sophisticated PLM architectures to serve as encoders for obtaining effective text representations.
-For example, Liu et al.
-Liu et al.
-(2022) proposed a dual-channel framework by modeling both literal and implied sentiments separately.
-They also constructed two conflict prompts to elicit PLMs to generate the sarcasm polarity Liu et al.
-(2023b).
-Qiao et al.
-Qiao et al.
-(2023) presented a mutual-enhanced incongruity learning network to take advantage of the underlying consistency between the two modules to boost the performance.
-Tian et al.
-Tian et al.
-(2023) proposed a dynamic routing Transformer network to activate different routing transformer modules for modeling the dynamic mechanism in sarcasm detection.
-However, the above-mentioned works still focus on how to utilize PLMs to extract effective features, without leveraging the extraodinary context learning capabilities of LLMs.
-In contrast, this paper makes the first attempt to explore the potential of prompting LLMs in sarcasm detection.
+(2022), etc., have been predominantly utilized during the middle stage of sarcasm detection research.
+Now, sarcasm detection research has stepped into the era of pre-trained language models (PLMs).
+An increasing number of researchers are designing sophisticated PLM architectures to serve as encoders for obtaining effective text representations Liu et al.
+(2023).
+Different from them, we propose four prompting methods to make the first attempt to explore the potential of prompting LLMs in sarcasm detection.
 3 The Proposed Framework: SarcasmCue
-The overall schematic illustration of the proposed SarcasmCue framework is illustrated in Fig.
+The proposed SarcasmCue framework is illustrated in Fig.
 2.
-We qualitatively compare SarcasmCue to other prompting approaches in Tab.
+We qualitatively compare SarcasmCue with other prompting approaches in Tab.
 1.
@@
 Scheme Seq? Non-Seq?
-Chain? Tree? Grap? Set? Tensor?
+Chain? Tree? Graph? Set? Tensor?
 […]
@@
 3.1 Task Definition
-Consider a sarcasm detection task.
 Given the data set $\mathcal{D}=\left\{\left(\mathcal{X},\mathcal{Y}\right)\right\}$, where $\mathcal{X}=\{x_{1},x_{2},\ldots,x_{n}\}$ denotes the input text sequence and $\mathcal{Y}=\{y_{1},y_{2},\ldots,y_{n}\}$ denotes the output label sequence.
 We use $\mathcal{L}_{\theta}$ to represent a large language model with parameter $\theta$.
-Our task is to leverage a collection of cues $\mathcal{C}=\{c_{1},c_{2},...,c_{k}\}$ to bridge the input $\mathcal{X}$ and the output $\mathcal{Y}$, where each cue $c_{i}$ is a coherent language sequence related to linguistics, context, or emotion that serves as an intermediate indicator toward identifying sarcasm.
+Our task is to leverage a collection of cues $\mathcal{C}=\{c_{1},c_{2},...,c_{k}\}$ to brige the input $\mathcal{X}$ and the output $\mathcal{Y}$, where each cue $c_{i}$ is a coherent language sequence that serves as an intermediate indicator toward identifying sarcasm.
 3.2 Chain of Contradiction
-We capture the inherent paradoxical nature of sarcasm, which is the incongruity between the surface sentiment and the true intention, and introduce chain of contradiction, a CoT-style paradigm that allows LLMs to decompose the problem of sarcasm detection into intermediate steps and solve each before making decision (Fig.
+We capture the inherent paradoxical nature of sarcasm, which is the incongruity between the surface sentiment and the true intention, and propose chain of contradiction, a CoT-style paradigm that allows LLMs to decompose the problem of sarcasm detection into intermediate steps and solve each before making decision (Fig.
 2 (a)).
@@
 Given the input sentence [$\mathcal{X}$], what is the SURFACE sentiment, as indicated by clues such as keywords, sentimental phrases, emojis?
-The output sequence $y_{1}\sim\mathcal{L}_{\theta}^{CoC}\left(\mathcal{Y}|p_{1}\right)$ is generated from the language model $\mathcal{L}_{\theta}^{CoC}$ conditioned on input prompt $p_{1}$.
-Step 2. We then ask LLM to carefully discover the true intention via the following prompt $p_{2}$:
-Deduce what the sentence really means, namely the TRUE intention, by carefully checking any rhetorical devices, language style, unusual punctuation, common senses.
-The output sequence, denoted as $y_{2}$, is generated from the language model conditioned on prompt $p_{2}$ as well as the previous interaction $p_{1},y_{1}$, formulated as $y_{2}\sim\mathcal{L}_{\theta}^{CoC}\left(\mathcal{Y}|p_{1},y_{1},p_{2}\right).$
-Step 3. We finally ask LLM to examine the consistency between surface sentiment and true intention and make the final prediction:
+$c_{1}$ is the output sequence, which can be formulated as $c_{1}\sim\mathcal{L}_{\theta}^{CoC}\left(c_{1}|\mathcal{X},p_{1}\right)$.
+Step 2. We thus ask LLM to carefully discover the true intention via the following prompt $p_{2}$:
+Deduce what the sentence really means, namely the TRUE intention, by carefully checking any rhetorical devices, language style, unusual punctuations, common senses.
+$c_{2}$ is the output sequence, which can be formulated as $c_{2}\sim\mathcal{L}_{\theta}^{CoC}\left(c_{2}|\mathcal{X},c_{1},p_{2}\right)$.
+Step 3. Let LLM examine the consistency between surface sentiment and true intention and make the final prediction:
 Based on Step 1 and Step 2, evaluate whether the surface sentiment aligns with the true intention.
@@
 Return the label only.
-$y_{3}$ is therefore generated based on a joint understanding of the preceding context $y_{1}$, $y_{2}$ and $p_{1}$, $p_{2}$, $p_{3}$: $y_{3}\sim\mathcal{L}_{\theta}^{CoC}\left(\mathcal{Y}|p_{1},y_{1},p_{2},y_{2},p_{3}\right)$.
-The sarcasm label is identified from $y_{3}$ as the output of CoC.
-Notably, CoC is built based on the presumption that all the cues are linearly correlated, and detects human sarcasm through step-by-step reasoning.
-Different from the original CoT, however, the steps are explicitly designed for the sarcasm detection context.
-Further details are presented in Algorithm 1 in App.
-A.
+CoC raises a presumption that the cues are linearly correlated, and detects human sarcasm through step-by-step reasoning. Further details see Algorithm 1 in App. A.
 3.3 Graph of Cues
 The linear structure of CoC restricts it to a single path of reasoning.
-To fill this gap, we introduce graph of cues, a GoT-style paradigm that allows LLMs to flexibly choose and weigh multiple cues, unconstrained by the need for unique predecessor nodes (Fig.
+To fill this gap, we introduce graph of cues, a graph based paradigm that allows LLMs to flexibly choose and weigh multiple cues, unconstrained by the need for unique predecessor nodes (Fig.
 2 (b)).
-GoC frames the problem of sarcasm detection as a search over a graph, and is formulated as a tuple $\left(\mathcal{M},\mathcal{G},\mathcal{E}\right)$, where $\mathcal{M}$ is the cue maker used to define what are the common cues, $\mathcal{G}$ is a graph of “sarcasm detection process”, $\mathcal{E}$ is cue evaluator used to determine which cues to keep selecting and in which order.
-Unlike ToT and GoT, GoC does not involve the modules of “thought generator” and “thought aggregation”.
+GoC frames the problem of sarcasm detection as a search over a graph, and is formulated as a tuple $\left(\mathcal{M},\mathcal{G},\mathcal{E}\right)$, where $\mathcal{M}$ is the cue maker used to define what are the common cues, $\mathcal{G}$ is a graph of “sarcasm detection process”, $\mathcal{E}$ is cue evaluator used to determine which cues to keep selecting.
 1.
@@
 Contextual cues refer to the environment and background of the text, including topic, cultural background, common knowledge.
-Emotional cues denote the emotional stance conveyed by the text, including emotional words, special symbols (such as emojis) and emotional contrasts.
-A total number of 4+3+3=10 cues are adopted.
+Emotional cues denote the emotions implied in the text, including emotional words, special symbols (such as emojis) and emotional contrasts.
+Hence, GoC can obtain 4+3+3=10 cues.
 2.
 Graph construction.
-In $\mathcal{G}=\left(V,E\right)$, the cues are regarded as vertices constituting the vertex set $V$, while the relations across cues form the edge set $E$.
-If there is an edge between cues $c_{k}$ and $c_{j}$, it is considered that $c_{k}$ and $c_{j}$ are closely related.
+In $\mathcal{G}=\left(V,E\right)$, 10 cues are regarded as vertices, constituting the vertex set $V$, the supplement relations across cues are regarded as edges.
 Given the cue $c_{k}$, the cue evaluator $\mathcal{E}$ considers cue $c_{j}$ to provide the most complementary information to $c_{k}$, which would combine with $c_{k}$ to facilitate a deep understanding of sarcasm.
@@
 Cue evaluator.
-We involve $\mathcal{G}$ in the LLM detecting sarcasm process.
-To advance this process, the cue evaluator $\mathcal{E}$ assesses the current progress towards judging sarcasm by means of determining whether the cumulative cues obtained so far are sufficient to yield an accurate judgment.
-If so, the search goes to an end.
-Otherwise, it serves as a heuristic for the search algorithm, determining which additional cues to select and in what order, to further the detection process.
-Similar to ToT, an LLM is used as the cue evaluator $\mathcal{E}$.
-We employ a voting strategy to determine the most valuable cue for selection, by explicitly comparing multiple potential cue candidates in a voting prompt, such as:
+We associate $\mathcal{G}$ with LLM detecting sarcasm process.
+To advance this process, the cue evaluator $\mathcal{E}$ assesses the current progress by asking the LLM whether the cumulative cues obtained thus far are sufficient to yield an accurate judgment.
+The search goes to an end if a positive answer is returned; otherwise, the detection process proceeds by instructing the LLM to determine which additional cues to select and in what order.
+In this work, an LLM will act as the cue evaluator, similar to ToT.
+We employ a voting strategy to determine the most valuable cue for selection, by deliberately comparing multiple potential cue candidates in a voting prompt, such as:
 Given an input text $\mathcal{X}$, the target is to accurately detect sarcasm.
@@
 This step can be formulated as $\mathcal{E}\left(\mathcal{L}_{\theta}^{GoC},c_{j+1}\right)\sim Vote\left\{\mathcal{L}_{\theta}^{GoC}\left(c_{j+1}|\mathcal{X},c_{1,2,...,j}\right)\right\}_{c_{j+1}\in\{c_{j+1},...,c_{k}\}}$.
-In a nutshell, it greedily selects the most valuable cue until the final judgment is reached.
-Although the GoC enables the exploration of many possible paths across the cue graph, its nature remains grounded in a step-by-step reasoning paradigm (see Algorithm 2 in App. A).
+Until the final judgment is reached, the most valuable cue are always selected in a greedy fashion.
+Although GoC enables the exploration of many possible paths across the cue graph, its nature remains grounded in a step-by-step reasoning paradigm (see Algorithm 2 in App.
+A).
 3.4 Bagging of Cues
-We further relax the assumption that the cues for sarcasm detection are inter-related.
-We introduce bagging of cues, an ensemble learning based paradigm that allows LLMs to independently consider varied combinations of cues without assuming a fixed order or dependency among them (Fig.
+We relax the assumption that the cues are interrelated in detecting sarcasm.
+We introduce bagging of cues, a ensemble learning based paradigm that allows LLMs to independently consider varied combinations of cues without assuming a fixed order or dependency among them (Fig.
 2 (c)).
-BoC constructs a pool of the pre-defined $k=10$ cues $\mathcal{C}$.
-From this pool, $\mathcal{T}$ subsets are random sampled, each consisting of $q~{}\left(i.e.,1\leq q\leq k\right)$ cues.
+BoC constructs a pool of the pre-defined 10 cues $\mathcal{C}$.
+From this pool, $\mathcal{T}$ subsets are obtained through $\mathcal{T}$ random samplings, where each subset $\mathcal{S}_{t}$ consists of $q~{}\left(i.e.,1\leq q\leq 10\right)$ cues.
 BoC thus leverages LLMs to generate $\mathcal{T}$ independent sarcasm predictions $\hat{y}_{t}$ based on the cues of each subset.
-Finally, such predictions are aggregated using a majority voting mechanism to produce the final sarcasm detection result.
-This approach embraces randomness in cue selection, enhancing the LLM’s ability to explore numerous potential paths, thus improving the robustness and accuracy of sarcasm detection.
-BoC consists of the following key steps:
+Finally, such predictions are aggregated using a majority voting mechanism to produce the final result.
+This approach embraces randomness in cue selection, enhancing the LLM’s ability to explore numerous potential paths.
+BoC consists of three key steps:
 Step 1.
 Cue subsets construction.
-A total of $\mathcal{T}$ cue subsets $\mathcal{S}_{t\in[1,2,...,\mathcal{T}]}=\left\{(c_{t_{1}},c_{t_{2}},...,c_{t_{q}}),t\in[1,2,...,\mathcal{T}]\right\}$ are created by randomly sampling without replacement from the complete pool of cues $\mathcal{C}$.
+A total of $\mathcal{T}$ cue subsets $\mathcal{S}_{t\in[1,2,...,\mathcal{T}]}=\left\{c_{t1},c_{t2},...,c_{tq}\right\}$ are created by randomly sampling without replacement from the complete pool of cues $\mathcal{C}$.
 Each sampling is independent.
@@
 LLM prediction.
-For each subset $\mathcal{S}_{t}$, an LLM $\mathcal{L}_{\theta}^{BoC}$ is used to independently make sarcasm prediction through the comprehensive analysis of the cues in the subset and the input text.
-This can be conceptually encapsulated as $\hat{y}_{t}\sim\mathcal{L}_{\theta}^{BoC}\left(\mathcal{Y}|\mathcal{S}_{t},\mathcal{X}\right)$.
-Step 3.
-Prediction aggregation.
-These individual predictions are then combined using an aggregation function, i.e., majority voting, to yield the final prediction: $Y\sim Vote\left(\left\{\hat{y}_{1},\hat{y}_{2},...,\hat{y}_{\mathcal{T}}\right\}\right)$.
-BoC treats all cues as independent and does not follow the step-by-step reasoning paradigm for sarcasm detection (see Algorithm 3 in App. A).
+For each subset $\mathcal{S}_{t}$, a LLM $\mathcal{L}_{\theta}^{BoC}$ is used to independently make sarcasm prediction through the comprehensive analysis of the cues in the subset and the input text.
+This can be conceptually encapsulated as $\hat{y}_{t}\sim\mathcal{L}_{\theta}^{BoC}\left(\hat{y}_{t}|\mathcal{S}_{t},\mathcal{X}\right)$.
+Step 3. Prediction aggregation. Such predictions $\{\hat{y}_{1},\hat{y}_{2},...,\hat{y}_{\mathcal{T}}\}$ are then combined using majority voting to yield the final prediction: $Y$.
+BoC does not follow the step-by-step reasoning paradigm for sarcasm detection (see Algorithm 3 in App. A.)
 3.5 Tensor of Cues
 CoC and GoC methods mainly handle low-order interactions between cues, while BoC assumes cues are independent.
-To capture high-order interactions among cues, we introduce tensor of cues, a novel paradigm that allows LLMs to amalgamate three types of cues ($viz.$ liguistic, contextual and emotional cues) into a high-dimensional representation (Fig.
+To capture high-order interactions among cues, we introduce tensor of cues, a stereo paradigm that allows LLMs to amalgamate three types of cues ($viz.$ linguistic, contextual and emotional cues) into a high-dimensional representation.
+(Fig.
 2 (d)).
@@
 We first ask the LLM to extract linguistic, contextual, and emotional cues respectively via a simple prompt.
-Taking linguistic cue extraction as an example:
-Instruction: Please extract the linguistic cues from the input sentence for sarcasm detection, such as keywords, rhetorical devices, punctuation and language style. Input: [$\mathcal{X}$]
+For example:
+Extract the linguistic cues from the input sentence for sarcasm detection, such as keywords, rhetorical devices, punctuation and language style.
 We take the outputs of the LLM’s final hidden layer as the embeddings of the linguistic, contextual and emotional cues, and apply a tensor fusion mechanism to fuse the cues as additional inputs to the sarcasm detection prompt.
@@
 (2017), we apply token-wise tensor fusion to aggregate the cues.
-In particular, the embeddings are projected on a low-dimensional space, i.e., $\vec{Lin}=\left(e_{1}^{l},e_{2}^{l},...,e_{L}^{l}\right)^{T}$, $\vec{Con}=\left(e_{1}^{c},e_{2}^{c},...,e_{L}^{c}\right)^{T}$, $\vec{Emo}=\left(e_{1}^{e},e_{2}^{e},...,e_{L}^{e}\right)^{T}$.
-Suppose the LLM has a hidden dimensionality of $d$, fully-connected layers $f_{lin},f_{con},f_{emo}$ are constructed to map the embeddings to dimensionality of $\{d_{l},d_{c},d_{e}\}$, respectively for linguistic, contextual and emotional cues.
+In particular, the embeddings are projected on a low-dimensional space via the fully-connected layers, i.e., $\vec{Lin}=\left(e_{1}^{l},e_{2}^{l},...,e_{L}^{l}\right)^{T}$, $\vec{Con}=\left(e_{1}^{c},e_{2}^{c},...,e_{L}^{c}\right)^{T}$, $\vec{Emo}=\left(e_{1}^{e},e_{2}^{e},...,e_{L}^{e}\right)^{T}$.
 Then, a tensor product is computed to combine the cues into a high-dimensional representation $\mathcal{Z}=\left(e_{1},e_{2},...,e_{L}\right)^{T}$, where
@@
 If you think the Input text is not sarcastic, answer: no.
-Input: [$\mathcal{X}$]
 The embedded prompt above is prepended with the aggregated cue sequence $\mathcal{Z}$ before fed to the LLM.
 As it is expected to output a single token of “yes” or “no” by design, we take the logit of the first generated token and decode the label accordingly as the output of ToC.
-ToC facilitates deep interactions among these cues, providing a powerful and flexible framework for processing complex linguistic phenomena (see Algorithm 4 in App.
+ToC facilitates deep interactions among these cues (see Algorithm 4 in App.
 A).
@@
 (2019).
-IAC-V1 and IAC-V2 are from the Internet Argument Corpus (IAC) Lukin and Walker (2013), specifically designed for the task of identifying and analyzing sarcastic remarks within online debates and discussions.
-It encompasses a balanced mixture of sarcastic and non-sarcastic comments.
-SemEval 2018 Task 3 is collected using irony-related hashtags (i.e.
-#irony, #sarcasm, #not) and are subsequently manually annotated to minimise the amount of noise in the corpuses.
-It emphasize the challenges inherent in identifying sarcasm within the constraints of MUStARD’s concise format, and highlight the importance of context and linguistic subtleties in recognizing sarcasm.
-MUStARD is compiled from popular TV shows including Friends, The Golden Girls, The Big Bang Theory, etc.
-It consists of 690 samples total of 3,000 utterances.
-Each sample is a conversation consisting of several utterances.
-In this work, we only use the textual information.
-The statistics for each dataset are shown in Table 2.
-Table 2: Dataset statistics.
-Dataset Avg. Length #Train #Dev #Test
-IAC-V1 68 1,595 80 320
-IAC-V2 43 5,216 262 1,042
-SemEval 2018 14 3,634 200 784
-MUStARD 14 552 - 138
+The details and statistics for each dataset are shown in Table 1 in App.
+B.
 Baselines. A wide range of SOTA baselines are included for comparison. They are:
-  • PLMs.
-(1) RoBERTa Liu et al.
-(2019), (2) BNS-Net Zhou et al.
-(2023), (3) DC-Net Liu et al.
-(2022), (4) QUIET Liu et al.
-(2023a) and (5) SarcPrompt Liu et al.
-(2023b) are five SOTA PLMs based approaches for sarcasm detection via pre-trained language modeling and refined representations.
-  • Prompt tuning.  (6) IO, (7) CoT Wei et al. (2022) and (8) ToT Yao et al. (2024) are four SOTA prompting approaches by leveraging advanced prompt approaches to enhance LLM’s performance.
-  • LLMs.  (9) GPT-4o [^2] https://openai.com/index/hello-gpt-4o/ and (10) LLAMA 3-8B-Instruct [^3] https://llama.meta.com/llama3/ are the strongest general LLMs.
+  • Prompt tuning.  (1) IO, (2) CoT Wei et al. (2022) and (3) ToT Yao et al. (2024) are three SOTA prompting approaches by leveraging advanced prompt approaches to enhance LLM’s performance.
+  • LLMs.
+We involve four general LLMs in the experiment, including (4) GPT-4o, (5) Claude 3.5 Sonnet, (6) Llama 3-8B and (7) Qwen 2-7B Bai et al.
+(2023).
+The first two are non-open-source LLMs while the last two are open-source LLMs.
+All four LLMs are representative of the strongest capabilities of their kinds.
 Implementation.
-We have implemented the prompting methods for GPT-4o and LLaMA3-8B-Instruct, and reported the performance of PLMs in their original papers.
-The GPT-4o methods are implemented with the official openAI Python API library [^4] https://github.com/openai/openai-python, while the LLaMA methods are implemented based on the Hugging Face Transformers library [^5] https://huggingface.co/docs/transformers.
-All prompting strategies are implemented for GPT-4o and LLaMA3-8B-Instruct except for ToC, which can solely be deployed on open-sourced LLMs.
-Following previous works in this field, LangChain [^6] https://github.com/langchain-ai/langchain is employed for the implementation of ToT and GoC.
-For the training of ToC, cross-entropy loss between the output logit and the true label token is computed to update the weights of the fully-connected layers.
-Table 3: Performance on four datasets. For LLMs, all strategies but ToC are based on a zero-shot setting.
+We have implemented the prompting methods for GPT-4o, Claude 3.5 Sonnet, Llama 3-8B and Qwen2-7B.
+The GPT-4o and Claude 3.5 Sonnet methods are implemented with the respective official Python API library: openAI [^3] https://github.com/openai/openai-python and anthropic [^4] https://github.com/anthropics/anthropic-sdk-python, while the LLaMA and Qwen methods are implemented based on the Hugging Face Transformers library [^5] https://huggingface.co/docs/transformers.
+Further details are presented in App.
+C.
+Table 2: Performance on four datasets.
+For LLMs, all strategies are based on a zero-shot setting.
+Blue and purple indicate the best and second-best results for each dataset.
+$\clubsuit$ represents significance improvement over the best baseline via unpaired t-test (p $<$ 0.05).
 Paradigm Method IAC-V1 IAC-V2 SemEval 2018 MUStARD Avg. of F1
 Acc. Ma-F1 Acc. Ma-F1 Acc. Ma-F1 Acc. Ma-F1
-PLMs RoBERTa 72.10 72.90 82.70 82.70 73.90 72.80 66.27 65.16 73.39
-BNS-Net 66.13 65.95 75.93 75.92 73.47 73.44 - - 71.77
-DC-Net 66.50 66.40 82.10 82.10 76.70 76.30 - - 74.93
-QUIET - - - - - - 72.36 72.13 -
-SarcPrompt 75.20 75.20 84.90 84.90 76.90 76.60 66.58 66.63 75.78
-GPT-4o IO 70.63 70.05 73.03 71.99 64.03 63.17 67.24 65.79 68.14
+GPT-4o IO 70.63 70.05 73.03 71.99 64.03 63.17 67.24 65.79 67.75
 CoT 61.56 58.49 58.83 56.42 58.92 51.99 58.11 55.76 55.67
-ToT 71.56 71.17 70.63 69.07 63.90 63.02 69.00 68.27 67.46
+ToT 71.56 71.17 70.63 69.07 63.90 63.02 69.00 68.27 67.88
 […]
-GoC (Ours) 85.38 68.08 64.97 61.30 74.03 74.02 70.69 69.91 68.33
-BoC (Ours) 68.75 67.36 71.35 69.39 62.12 61.85 69.42 68.45 66.79
-LLaMA 3-8B-Instruct IO 55.94 46.40 54.70 43.74 49.36 44.46 54.64 44.99 44.90
+GoC (Ours) 65.00 62.91 64.97 61.30 74.03 74.02 70.69 69.91 67.04
+BoC (Ours) 68.75 67.36 71.35 69.39 62.12 61.85 69.42 68.45 66.76
+Claude 3.5 Sonnet IO 66.56 66.54 76.78 76.62 75.13 75.11 74.78 74.78 73.26
+CoT 71.25 71.14 74.66 74.10 71.56 71.47 73.62 73.53 72.56
+ToT 63.44 62.48 71.88 71.74 68.62 68.61 58.84 54.46 64.32
+[…]
+GoC (Ours) 70.94 70.93 74.67 74.18 76.91 76.91 70.00 69.85 72.97
+BoC (Ours) 66.88 66.40 73.61 72.82 70.28 70.07 72.61 71.93 70.31
+Llama 3-8B IO 55.94 46.40 54.70 43.74 49.36 44.46 54.64 44.99 44.90
 CoT 56.25 47.28 54.22 42.96 49.36 44.55 54.20 44.86 44.91
@@
 […]
-GoC (Ours) 57.10 54.96 42.20 41.61 57.33 57.24 52.77 52.67 51.62
-BoC (Ours) 62.50 59.28 62.57 58.11 59.82 58.40 59.71 56.70 58.12
-ToC (Ours) 70.31 70.29 79.08 79.07 77.93 76.86 73.33 72.85 74.77
+GoC (Ours) 57.10 54.96 54.22 53.30 57.33 57.24 52.77 52.67 54.54
+BoC (Ours) 62.50 59.28 62.57 58.11 65.94 65.50 59.71 56.70 59.90
+ToC (Ours) 62.19 61.78 72.95 72.94 68.88 68.21 61.26 58.03 65.24
+Qwen 2-7B IO 56.56 49.32 51.82 38.57 45.15 38.83 54.78 46.17 43.22
+CoT 54.69 46.53 52.88 40.12 43.24 35.79 54.93 45.81 42.06
+ToT 53.44 43.71 50.29 39.62 44.26 38.12 52.90 44.60 41.51
+[…]
+GoC (Ours) 55.00 47.35 53.45 42.25 45.03 38.17 54.49 47.49 43.82
+BoC (Ours) 52.50 43.78 52.40 40.24 49.87 45.63 54.06 46.11 43.94
+ToC (Ours) 71.56 71.56 72.33 71.76 68.88 68.77 65.94 61.46 68.39
 4.2 Main Results
-We report both Accuracy and Macro-F1 results for SarcasmCue and baselines in a zero-shot setting in Table 3, except for ToC which requires supervised training for context adaption.
-LLMs do not possess a unique advantage on sarcasm detection.
-Since sarcasm indicates the manifestation of sentiments and intentions opposite to the literal meaning of the texts, it usually violates logical reasoning pipelines that LLMs are known to excel at Wei et al.
+We report both Accuracy and Macro-F1 scores for SarcasmCue and baselines in Table 2.
+(1) SarcasmCue consistently outperforms SoTA prompting baselines.
+The proposed prompting strategies in the SarcasmCue framework achieve an overall superior performance compared to the baselines and consistently push the SoTA by 4.2%, 2.0%, 29.7% and 58.2% on F1 scores across four datasets.
+In particular, by explicitly designing the reasoning steps for sarcasm detection, CoC beats CoT by a tremendous margin on GPT-4o and Claude 3.5 Sonnet, whilst performing in par with CoT on Llama 3-8B and Qwen 2-7B.
+By pre-defining the set of cues in three main categories, GoC and BoC effectively guide LLMs to reason along correct paths, leading to more accurate judgments of sarcasm compared to the freestyle thinking in ToT.
+For example, the best proposed method, CoC (74.74), brings a 2.0% improvement over the best baseline method, IO (73.26).
+ToC achieves an effective tensor fusion of multi-aspect cues for sarcasm detection, significantly outperforming other baselines.
+For instance, it exhibits a 29.7% improvement over the best baseline method, ToT (50.31).
+(2) Sarcasm detection does not necessarily follow a step-by-step reasoning process.
+The comparison between sequential (CoT, CoC, GoC, ToT) and non-sequential (BoC, ToC) prompting strategies fails to provide clear empirical evidences on whether sarcasm detection follows a step-by-step reasoning process.
+Nevertheless, the results on Llama 3-8B are more indicative to GPT-4o and Claude 3.5 Sonnet, since the latter models have strong capabilities on their own (IO) and do not significantly benefit from any prompting strategies.
+For Llama 3-8B and Qwen 2-7B, non-sequential methods, particularly ToC, show superior performance.
+In Llama 3-8B, ToC achieves an average F1 score of 65.24%, which is 8.9% higher than the best sequential method (GoC at 54.54%).
+The difference is even more pronounced on Qwen 2-7B.
+This seems to support our hypothesize that sarcasm has a non-sequential nature.
+Table 3: Ablation study of BoC, GoC and ToC. All strategies are run on a zero-shot setting. The best results for each dataset are colored in blue.
+LLMs Method IAC-V1 IAC-V2 SemEval MUStARD Avg. of F1
+Claude 3.5 w/o Lin 68.41 75.62 77.42 69.66 72.78
+w/o Emo 69.65 74.04 78.70 70.57 73.24
+w/o Con 70.53 74.91 76.39 70.11 72.99
+GoC 70.93 74.18 76.91 69.85 72.97
+w/o Lin 45.89 42.49 47.47 65.33 50.30
+w/o Emo 58.00 56.99 56.81 68.84 60.16
+w/o Con 61.71 63.70 69.53 74.80 67.44
+BoC 66.40 72.82 70.07 71.93 70.31
+Llama 3-8B w/o Lin 45.79 51.90 56.01 46.84 50.14
+w/o Emo 48.60 49.40 52.38 45.12 48.88
+w/o Con 52.51 53.69 52.14 48.28 51.66
+GoC 54.96 53.30 57.24 52.67 54.54
+w/o Lin 52.71 57.51 57.53 53.06 55.20
+w/o Emo 57.33 59.40 62.01 53.06 57.95
+w/o Con 56.88 60.36 59.04 52.30 57.15
+BoC 59.28 58.11 65.50 56.70 59.90
+w/o Lin 53.31 67.05 59.20 48.05 56.90
+w/o Emo 57.42 67.08 64.01 52.89 60.35
+w/o Con 55.26 71.78 63.93 52.48 60.86
+ToC 61.78 72.94 68.21 58.03 65.24
+4.3 Ablation Study
+Table 3 presents the result of ablation study.
+w/o Lin, w/o Emo, w/o Con refer to the method where linguistic, emotional and contextual cues are ablated, respectively.
+To avoid proactive extraction of ablated cues by an LLM, we explicitly “prompt away” the cues in the inputs.
+An example prompt could be “You can only use the emotional cues and contextual cues, and do not use any linguistic information here” for the w/o Lin case.
+The experiment results highlight the following conclusions: (a) the removal of any single type of cue leads to a noticeable drop in performance across all datasets, demonstrating the importance of each type of cue in sarcasm detection; (b) linguistic cues appear to have the most significant impact, as removing them leads to a noticeable decrease in performance across most settings; (c) the absence of contextual cues also affects the performance, but to a lesser extent compared to linguistic cues.
+4.4 Zero-shot v/s Few-shot Prompting
+Since the above experiments are mainly based on a zero-shot setting, we are curious of whether the conclusions also apply in a few-shot scenario.
+Therefore, we perform few-shot experiments to evaluate whether the proposed SarcasmCue framework can perform better when a limited number of contextual examples are available.
+We plot the main results in Fig.
+3, we randomly sample $k=\left\{0,1,5,10\right\}$ examples from the training set.
+Please refer to Table 2, App.
+D for the full result.
+Figure 3: The average Macro-F1 across K-shots for the GPT-4o and Claude 3.5 Sonnet models.
+As shown in the plot, the number of demonstrations has a significant impact on the results.
+For example, CoC appears sensitive to the initial introduction of demonstration examples with a slight descent in performance when only 1 example is provided.
+However, as the number of shots increases to 5 and 10, the performance progressively improves.
+This trend underscores the effectiveness of CoC in adapting and refining its approach with more examples.
+In contrast, BoC demonstrates a consistent improvement in performance as the number of shots increases.
+Overall, these results demonstrate the robustness and adaptability of the SarcasmCue framework in zero-shot and few-shot scenarios.
+The framework can effectively utilize limited contextual examples to further improve sarcasm detection, making it suitable for applications where large annotated datasets are not readily available.
+Figure 4: The influence of model scale. The figures in the top and bottom correspond to Qwen and Llama models, respectively.
+4.5 Influences of LLM scales
+In an attempt to study the influence of different LLM scales, we evaluate the performance of sarcasm detection of Qwen and Llama of varying sizes, see Fig. 4.
+The key take-aways are two-fold.
+First, the efficacy of our prompting methods is amplified with increasing model scale.
+This aligns closely with the key findings of the CoT method Wei et al.
 (2022).
-This is empirically validated in the experiment where LLMs are observed to have consistently lower performance over PLMs in terms of average F1 scores across the four datasets.
-This highlights the need to investigate prompting strategies for adapting LLMs for sarcasm detection, towards which this work has made the first attempt and achieved preliminary success.
-Human sarcasm detection does not necessarily follow a step-by-step reasoning process.
-The comparison between sequential (CoT, CoC, GoC, ToT) and non-sequential (BoC, ToC) prompting strategies fails to provide clear empirical evidences on whether sarcasm detection follows a step-by-step reasoning process.
-Nevertheless, the results on LLaMA3-8B-Instruct are more indicative to GPT-4o, since the latter has a strong capacity on its own (IO) and does not significantly benefit from any prompting strategies on its top.
-On LLaMA3-8B-Instruct where in-context learning is necessary for sarcasm detection due to its poor IO performance, non-sequential approaches can apparently offer more benefits over sequential ones, with a remarkable margin consistently present on all four datasets.
-This seems to support our hypothesize that sarcasm has a non-sequential nature.
-SarcasmCue successfully adapts LLMs to sarcasm detection.
-The proposed prompting strategies in the SarcasmCue framework achieve an overall superior performance to the baseline prompting methods and bring about accuracy increase over the original LLMs in a zero-shot setting.
-In particular, by explicitly designing the reasoning steps for sarcasm detection, CoC beats CoT by a tremendous margin on GPT-4o, whilst performing in par with CoT on LLaMA3-8B-Instruct, an interesting result that further suggests the non-squential nature of sarcasm detection.
-By pre-defining the set of cues on 3 main aspects, GoC and BoC manage to guide LLMs to reason along the correct paths, leading to more accuracy judgment of sarcasm than the freestyle thinking in ToT.
-The proposed trainable neural architecture in ToC achieves an effective tensor fusion of multi-aspect cues for sarcasm detection, pushing the capacity to a comparable level to PLMs without tuning the LLM parameters.
-5 Acknowledgments
-This work is supported by National Science Foundation of China under grant No.
-62006212, Fellowship from the China Postdoctoral Science Foundation (2023M733907), Natural Science Foundation of Hunan Province of China (242300421412).
-6 Conclusion
-In this work, we aim to study the step-wise reasoning nature of sarcasm detection, and introduce a prompting framework (called SarcasmCue) containing four sub-methods, $viz.$ chain of contradiction (CoC), graph of cues (GoC), bagging of cues (BoC) and tensor of cues (ToC).
-It elicits LLMs for human sarcasm detection by considering sequential and non-sequential prompting methods.
-Our comprehensive evaluations across multiple benchmarks and state-of-the-art LLMs demonstrate that SarcasmCue outperforms traditional methods, with non-sequential prompting methods (GoC and ToC) showing particularly strong performance.
-In the future, we plan to develop the multi-modal version of SarcasmCue for multi-modal sarcasm detection.
-7 Limitations
-The proposed SarcasmCue model has several limitations: (1) It incorporates only three types of cues – linguistic, contextual, and emotional – while other potentially useful cues, such as multimodal information, have not been integrated, potentially limiting the model’s comprehensive understanding of sarcasm; (2) the performance of SarcasmCue is influenced by the capabilities of the underlying large language models (LLMs), meaning it performs better with more powerful LLMs.
-Appendix A Algorithms of Four Prompting Methods
+This occurs because when an LLM is sufficiently large, its capabilities for multi-hop reasoning and understanding language are significantly enhanced.
+Second, ToC exhibits high sensitivity to model scale, performing significantly better in larger models, making it particularly suitable for larger-scale applications.
+CoC and GoC demonstrate moderate sensitivity, indicating a balance between performance improvement and scalability.
+BoC offers robust performance even in smaller models, suggesting its utility in resource-constrained scenarios.
+Overall, our proposed framework has a high adaptability across various model scales by offering suitable methods.
+Please see Table 3 and Fig.
+1, App.
+E for the full results.
+4.6 Error Analysis
+Figure 5: The average error rate of the four prompting methods.
+Fig.
+5 shows the error rates of failure cases in terms of false negative (FN) and false positive (FP) for all four prompting methods in SarcasmCue.
+CoC, GoC and BoC exhibit higher false positive rates, indicating an over-detection of sarcasm that could lead to the frequent misclassification of normal statements as sarcastic.
+In contrast, ToC exhibits the lowest overall error rate and the FP and FN rates are indeed much closer to each other, indicating a balanced performance in detecting both sarcastic and non-sarcastic texts.
+These insights highlight potential directions for future improvements in sarcasm detection methodologies.
+The higher false positive rates suggest a need for refining these methods to reduce over-sensitivity and improve discrimination between sarcastic and non-sarcastic texts.
+The detailed case study is presented in App.
+F.
+4.7 Extension to New Task
+To evaluation the generalization capability of SarcasmCue, we apply it to another complex affection understanding task, humor detection.
+We compare our proposed SarcasmCue (where the backbone is GPT-4o) with two supervised PLMs (MFN Hasan et al.
+(2021) and SVM+BERT Zhang et al.
+(2024)) on two benchmarking datasets, CMMA Zhang et al.
+(2024) and UR-FUNNY-V2 Hasan et al.
+(2019).
+As shown in Table 4, our methods (BoC and CoC) surpass the baseline on CMMA, whilst performing in par to the strongest baselines on the UR-FUNNY-V2 dataset.
+These results highlight the strong generalizability and versatility of our framework, confirming its potential utility across a wide range of affection understanding tasks.
+Table 4: Performance on two humor detection datasets.
+Method CMMA UR-FUNNY-V2 Avg. of F1
+Acc. Ma-F1 Acc. Ma-F1
+MFN - - 64.44 64.12 -
+SVM+BERT 55.23 54.08 69.62 69.27 61.68
+CoC 78.14 58.60 64.08 60.13 65.24
+GoC 79.60 57.42 64.89 61.65 65.89
+BoC 75.81 58.58 68.71 66.83 67.48
+5 Conclusions
+This work aims to study the stepwise reasoning nature of sarcasm detection, and introduces a prompting framework (called SarcasmCue) containing four sub-methods, $viz.$ CoC, GoC, BoC and ToC.
+It elicits LLMs to detect human sarcasm by considering sequential and non-sequential prompting methods.
+Our comprehensive evaluations across multiple benchmarks and SoTA LLMs demonstrate that SarcasmCue outperforms traditional methods and pushes the state-of-the-art by 4.2%, 2.0%, 29.7% and 58.2% F1 scores across four datasets.
+Additionally, the performance of SarcasmCue on humor detection further validate its robustness and versatility.
+Limitations.
+SarcasmCue has its limitation: it incorporates only three types of cues, while other potentially useful cues have not been integrated, potentially limiting the model’s comprehensive understanding of sarcasm.
+Appendix A A. Algorithms of Four Prompting Methods
 1. CoC. We present further details of CoC in Algorithm 1.
@@
 16:return $\mathcal{Y}$
+3. BoC. We present further details of BoC in Algorithm 3.
 Algorithm 3 Bagging of cues
@@
 14:return $Y$
-3. BoC. We present further details of BoC in Algorithm 3.
 4. ToC. We present further details of ToC in Algorithm 4.
@@
 7:Step 2: Construct Tensor Representation
-8:Compute tensor product to combine cues: $\mathcal{Z}=\vec{Lin}\otimes\vec{Con}\otimes\vec{Emo}$
+8:Compute tensor product to combine cues: $\mathcal{Z}=\begin{bmatrix}\vec{Lin}\\ 1\end{bmatrix}\otimes\begin{bmatrix}\vec{Con}\\ 1\end{bmatrix}\otimes\begin{bmatrix}\vec{Emo}\\ 1\end{bmatrix}$
 9:Step 3: Sarcasm Detection
@@
 12:return $\mathcal{Y}$
+Appendix B B. Datasets Details
+Datasets.
+Four benchmarking datasets are selected as the experimental beds, $viz.$ IAC-V1 Lukin and Walker (2013), IAC-V2 Oraby et al.
+(2016), SemEval 2018 Task 3 Van Hee et al.
+(2018) and MUStARD Castro et al.
+(2019).
+Table 5: Dataset statistics.
+Dataset Avg. Length #Train #Dev #Test
+IAC-V1 68 1,595 80 320
+IAC-V2 43 5,216 262 1,042
+SemEval 2018 14 3,634 200 784
+MUStARD 14 552 - 138
+IAC-V1 and IAC-V2 are from the Internet Argument Corpus (IAC) Lukin and Walker (2013), specifically designed for the task of identifying and analyzing sarcastic remarks within online debates and discussions.
+It encompasses a balanced mixture of sarcastic and non-sarcastic comments.
+SemEval 2018 Task 3 is collected using irony-related hashtags (i.e.
+#irony, #sarcasm, #not) and are subsequently manually annotated to minimise the amount of noise in the corpuses.
+It emphasize the challenges inherent in identifying sarcasm within the constraints of MUStARD’s concise format, and highlight the importance of context and linguistic subtleties in recognizing sarcasm.
+MUStARD is compiled from popular TV shows including Friends, The Golden Girls, The Big Bang Theory, etc.
+It consists of 690 samples total of 3,000 utterances.
+Each sample is a conversation consisting of several utterances.
+In this work, we only use the textual information.
+The statistics for each dataset are shown in Table 5.
+Appendix C C. Implementation Details
+We have implemented the prompting methods for GPT-4o, Claude 3.5 Sonnet, LLaMA3-8B-Instruct and Qwen 2-7B.
+The GPT-4o and Claude 3.5 Sonnet methods are implemented with the respective official Python API library: openAI [^6] https://github.com/openai/openai-python and anthropic [^7] https://github.com/anthropics/anthropic-sdk-python, while the LLaMA and Qwen methods are implemented based on the Hugging Face Transformers library [^8] https://huggingface.co/docs/transformers.
+All prompting strategies are implemented for GPT-4o and Claude 3.5 Sonnet except for ToC, which can solely be deployed on open-sourced LLMs.
+Following previous works in this field, LangChain [^9] https://github.com/langchain-ai/langchain is employed for the implementation of ToT and GoC.
+For the training of ToC, cross-entropy loss between the output logit and the true label token is computed to update the weights of the fully-connected layers.
+The mean performance of each model over 5 runs is calculated.
+Given the proprietary nature of GPT-4o and Claude 3.5 Sonnet, we have implemented only CoC, GoC and BoC prompting approaches.
+For Llama 3-8B and Qwen 2-7B, we implemented all four proposed prompting approaches.
+This is due to the reasons previously discussed: ToC requires access to and modification of the base model.
+We run all the models on four A100 GPUs.
+Table 6: Few shot performance testing.
+LLMs Method K-shot IAC-V1 IAC-V2 SemEval MUStARD Avg. of F1
+GPT-4o CoC 0-shot 71.52 72.31 70.60 68.48 70.73
+1-shot 58.91 61.99 63.69 60.98 61.39
+5-shot 60.65 66.00 65.39 63.60 63.91
+10-shot 63.34 70.29 67.59 63.40 66.16
+GoC 0-shot 62.91 61.30 74.02 69.91 67.04
+1-shot 66.00 66.70 73.00 64.12 67.45
+5-shot 66.93 65.88 73.41 67.77 68.50
+10-shot 74.38 66.36 69.34 68.61 69.67
+BoC 0-shot 67.36 69.39 61.85 68.45 66.76
+1-shot 64.67 66.36 61.08 73.47 66.40
+5-shot 65.66 67.52 64.06 73.97 67.80
+10-shot 70.70 69.06 68.12 76.51 71.10
+Claude 3.5 Sonnet CoC 0-shot 69.40 73.17 82.23 74.16 74.74
+1-shot 72.45 74.38 77.29 72.75 74.22
+5-shot 71.99 78.12 79.42 75.65 76.30
+10-shot 75.49 79.07 83.46 79.56 79.40
+GoC 0-shot 70.93 74.18 76.91 69.85 72.97
+1-shot 65.16 69.29 74.98 65.40 68.71
+5-shot 69.01 72.76 75.62 67.12 71.13
+10-shot 72.80 74.71 76.65 68.36 73.13
+BoC 0-shot 66.40 72.82 70.07 71.93 70.31
+1-shot 74.63 81.09 76.64 75.38 76.94
+5-shot 78.40 84.34 79.72 82.83 81.32
+10-shot 80.27 84.82 82.76 85.76 83.40
+Appendix D D. Zero-shot v/s Few-shot Prompting
+We perform zero-shot and few-shot experiments to evaluate whether the proposed SarcasmCue framework can perform better when a limited number of contextual examples are available.
+The results are shown in Table 6.
+We design four $k$-shot settings: zero-shot, one-shot, five-shot, ten-shot.
+For each setting, we randomly sample $k=\left\{0,1,5,10\right\}$ examples from the training set.
+The impact of adding shots varies with the number of shots.
+For example, CoC appears sensitive to the initial introduction of demonstration examples with a slight descent in performance when only 1 example is provided.
+However, as the number of shots increases to 5 and 10, the performance progressively improves.
+This trend underscores the effectiveness of CoC in adapting and refining its approach with more examples.
+In contrast, BoC demonstrates a consistent improvement in performance as the number of shots increases.
+Compared to CoC and BoC, GoC exhibits a relatively lower sensibility to the presence of demonstration examples, while still showing a slight but stable improvement with more shots.
+Overall, these results demonstrate the robustness and adaptability of the SarcasmCue framework in zero-shot and few-shot scenarios.
+The framework can effectively utilize limited contextual examples to improve sarcasm detection, making it suitable for applications where large annotated datasets are not readily available.
+This adaptability underscores the practical value of SarcasmCue in real-world settings where training data may be scarce.
+Table 7: Influence of model scale. Macro-F1 score is measured on all four datasets, and the average Macro-F1 score is computed and shown in the last column.
+LLMs Method IAC-V1 IAC-V2 SemEval MUStARD Avg. of F1
+Qwen 2-1.5B CoC 48.05 44.43 44.05 50.66 46.80
+GoC 43.75 53.21 50.69 45.63 48.32
+BoC 43.84 42.86 42.87 52.41 45.49
+ToC 57.46 57.60 60.69 54.40 57.53
+Qwen 2-7B CoC 45.77 38.90 36.37 44.26 41.33
+GoC 47.35 42.25 38.17 47.49 43.82
+BoC 43.78 40.24 45.63 46.11 43.94
+ToC 71.56 71.76 68.77 61.46 68.39
+Qwen 2-72B CoC 61.33 59.42 44.92 51.63 54.33
+GoC 57.67 68.78 65.28 61.87 63.40
+BoC 45.24 43.38 44.18 48.10 45.23
+LlaMA 3-8B CoC 46.95 42.60 44.36 45.66 44.89
+GoC 54.96 53.30 57.24 52.67 54.54
+BoC 59.28 58.11 65.50 56.70 59.90
+LlaMA 3-70B CoC 68.94 77.29 62.59 59.73 67.14
+GoC 56.83 62.57 58.66 53.81 57.97
+BoC 65.49 68.52 55.14 45.72 58.72
+Figure 6: The influence of model scale.
+Appendix E E. Influences of LLM scales
+In an attempt to study the influence of different LLM scales, we evaluate the performance of sarcasm detection of Qwen and Llama of varying sizes.
+Table 7 presents the macro-F1 scores of each model across the four sarcasm detection tasks.
+The key take-aways are two-fold.
+First, with increasing model scale, the efficacy of our prompting is exponentially amplified.
+This aligns closely with the key findings of the CoT method Wei et al.
+(2022).
+This is because when an LLM is sufficiently large, its capabilities for multi-hop reasoning are greatly developed and strengthened.
+More specifically:
+(1) CoC demonstrates a significant improvement in performance as model scale increases.
+For Qwen models, the average F1 score rises from 46.80% (1.5B) to 54.33% (72B).
+LLaMA models show an even more pronounced enhancement, with the average F1 score jumping from 44.89% (8B) to 67.14% (70B).
+This indicates that CoC becomes more effective with larger model scales.
+(2) GoC also exhibits a positive trend with increasing model size.
+In Qwen models, performance improves from 48.32% (1.5B) to 63.40% (72B) average F1 score.
+LLaMA models display a similar trend, with the average F1 score increasing from 54.54% (8B) to 57.97% (70B).
+These results suggest that GoC generally benefits from larger model scales across different architectures.
+(3) BoC shows inconsistent performance across model scales.
+For Qwen models, performance remains relatively stable, with a slight decrease in the 72B model (45.23% average F1) compared to smaller versions.
+LLaMA models demonstrate a minor decline in performance, with the average F1 score decreasing from 59.90% (8B) to 58.72% (70B).
+This suggests that BoC might be more effective with smaller model scales.
+(4) ToC exhibits the most substantial improvement within the available data range. For Qwen models, the average F1 score increases dramatically from 57.53% (1.5B) to 68.39% (7B).
+Overall, our proposed framework demonstrates high adaptability across different model scales by offering a range of methods.
+This adaptability allows for optimized performance based on available computational resources and specific task requirements
+Table 8: Typical examples for case study.
+Example Text Golden CoC GoC BoC ToC
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Appendix F F. Case Study
+We analyze the proposed four prompting approaches on several typical cases in Table 8.
+We categorize and analyze sarcasm detection methods.
+In scenarios involving straightforward statements (Examples 8, 15), all methods correctly identify texts as non-sarcastic, showcasing the SarcasmCue framework’s efficacy in clear-cut non-sarcastic contexts.
+For scenarios marked by clear linguistic contrasts (Examples 1, 6, 7), the CoC and GoC methods demonstrate superior performance.
+They effectively capture textual contradictions, making them ideally suited for texts where the apparent meaning sharply diverges from the intended message.
+For texts involving complex contexts that necessitate an understanding of nuanced background knowledge (Examples 2, 3, 9, 10), the BoC and ToC methods prove more effective.
+BoC achieves this through sampling multiple subsets of cues, thus capturing the complexity of the context, whereas ToC employs a multi-view representation to process intricate high-order interactions.
+In scenarios characterized by subtle sarcasm (Examples 5, 11)—where texts may lack overt sarcastic markers or structural clues—ToC outperforms other methods.
+It excels in capturing the intricate interaction among linguistic, contextual, and emotional cues.
+Additionally, for texts involving specialized domain knowledge (Examples 13, 14), both BoC and ToC are effective due to their ability to integrate and analyze domain-specific cues.
+This analysis highlights that different sarcasm detection methods are tailored to specific textual scenarios.
+CoC and GoC are highly effective in environments with straightforward linguistic oppositions, where the sarcasm is direct and easily discernible.
+Conversely, BoC and ToC are particularly adept in scenarios that demand a deeper understanding of complex and subtle cues.
+ToC is especially notable for its performance across a broad range of scenarios, attributed to its capability to capture and analyze complex interactions among multiple layers of cues.
+However, in highly ambiguous situations, a blend of methods or the addition of extra contextual information may be required.
+This insight directs future research towards identifying or combining the most appropriate methods for enhancing the overall accuracy of sarcasm detection across varied scenarios.